### v1

# On a magneto-spectral invariant on finite graphs

Abstract.
In this paper, we introduce a magneto-spectral invariant for finite graphs. This invariant vanishes on trees and is maximized by complete graphs. We compute this invariant for cycles, complete graphs, wheel graphs and suspensions of trees and derive various lower and upper bounds. We discuss the behaviour of this invariant under various graph operations and investigate relations to the spectral gap.
1. Introduction
In this paper, we introduce a new magneto-spectral invariant (see Definition 1.2 below) and discuss various of its properties. We begin by providing the notational background.
Let $G=(V,E)$ be a finite graph. For adjacent vertices $x,y\in V$ we use the notation $x\sim y$. The set of oriented edges $E^{or}(G)$ consists of ordered pairs $(x,y)$ with $x\sim y$. A magnetic potential is a map $\sigma:E^{or}(G)\to S^{1}$ with $\sigma(y,x)=\sigma(xy)^{-1}$. In the case that $\sigma$ assumes only the values ±$\pm 1$, we also refer to it as a signature. The magnetic Laplacian associated to a magnetic potential $\sigma$ is a linear operator $\Delta^{\sigma}:C(V,\mathbb{C})\to C(V,\mathbb{C})$, given by
(1) $\Delta^{\sigma}f(x)=\sum_{y\sim x}(f(x)-\sigma(x,y)f(y)).$
In the case of $\sigma\equiv 1$, we recover the standard non-normalized Laplacian $\Delta$. The space $C(V,\mathbb{C})$ has the following inner product: Let $f,g\in C(V,\mathbb{C})$. Then
$\langle f,g\rangle=\sum_{x\in V}f(x)\overline{g(x)}.$
$\Delta^{\sigma}$ is symmetric and has non-negative real spectrum. Its smallest eigenvalue is denoted by $\lambda_{1}^{\sigma}(G)$ and its variational characterization is given by
$\lambda_{1}^{\sigma}(G)=\inf_{f\neq 0}\mathcal{R}^{\sigma}(f)$
with
$\mathcal{R}^{\sigma}(f)=\frac{\sum_{\{x,y\}\in E}|f(x)-\sigma(x,y)f(y)|^{2}}{\sum_{x\in V}|f(x)|^{2}}.$
The eigenvalues of the standard Laplacian $\Delta$ on $G$ are denoted by
$0=\lambda_{1}(G)\leq\lambda_{2}(G)\leq\cdots\leq\lambda_{N}(G)$
with $N=|V|$. The spectral gap of a graph $G$ is defined as $\lambda_{2}(G)$.
An important transformation of magnetic potentials is given in the following definition.
Definition 1.1.
Let $G=(V,E)$ be a finite graph and $\sigma:E^{or}(G)\to S^{1}$. For any function $\tau:V\to S^{1}$, we define the associated gauge transformation $\sigma^{\tau}$ of $\sigma$ as
$\sigma^{\tau}(x,y)=\tau(x)^{-1}\sigma(x,y)\tau(y).$
The function $\tau$ is called a gauge function. We say that two magnetic potentials $\sigma_{1},\sigma_{2}$ are gauge equivalent, if they are related via a suitable gauge function.
Note that magnetic Laplacians with gauge-equivalent magnetic potentials are unitarily equivalent and have therefore the same spectrum. Note also that any magnetic potential can be gauged to be trivial on a given spanning forest.
In the case of signatures, we also refer to gauge functions $\tau:V\to\{-1,1\}$ as switching functions and refer to two signatures $\sigma_{1},\sigma_{2}$ which are related via a switching function as being switching equivalent. Signatures which are switching equivalent to the trivial signature are called balanced. We denote the trivial signature by $\sigma_{0}\equiv 1$ in this paper. The signature assuming only values $-1$ is referred to as the anti-balanced signature. It is denoted by $\overline{\sigma}\equiv-1$. We refer to magnetic Laplacians corresponding to signatures as signed Laplacians.
Signed graphs and signed Laplacians have been studied extensively (see, e.g. [AL20, H57, Z82, Z98, Z13] and references therein). On the other hand, there is a vast literature on discrete magnetic Laplacians: a small collection of some papers on this topic is [B13, BGKLM20, CdV13, DM06, FCLP18, G14, HSh01, KS23, LLPP15, Shu94, Su94].
Now we introduce the following magneto-spectral invariant of finite graphs.
Definition 1.2 (Magneto-spectral height).
Let $G=(V,E)$ be a finite graph. The magneto-spectral height of $G$ is defined to be
(2) $\nu(G):=\sup_{\sigma}\lambda_{1}^{\sigma}(G),$
where the supremum is taken over all magnetic potentials $\sigma:E^{or}(G)\to S^{1}$.
A magnetic potential is called maximal, if $\lambda_{1}^{\sigma}(G)$ agrees with $\nu(G)$.
Henceforth, all our graphs are assumed to be finite.
1.1. Structure of the paper
In Section 2, we compute the magneto-spectral height for various examples, namely, we have $\nu(T)=0$ for trees (Lemma 2.1), $\nu(C_{d})=2-2\cos(\pi/d)$ for cycles $C_{d}$ (Example 2.2), $\nu(K_{d})=d-2$ for complete graphs $K_{d}$ (Example 2.6), and $\nu(W_{d})=3-2\cos(\pi/d)$ for wheel graphs $W_{d}$ (Proposition 2.13). It is known (see [ZLY09]) that $W_{6}$ is not determined by its Laplace spectrum and that there exists a cospectral graph $\widehat{G}$. We show that the magneto-spectral height distinguishes $W_{6}$ and $\widehat{G}$ (Subsection 2.4). Moreover, we also discuss some counterexamples to natural questions in Subsection 2.5.
Section 3 is concerned with various upper and lower bounds for the magneto-spectral height. While the derivation of lower bounds is often very challenging, upper bounds can be derived more easily. We present combinatorial upper bounds (Proposition 3.1), upper bounds in terms of topological/spectral data of subgraphs (Theorem 3.3), upper and lower Ramanujan type bounds (Propositions 3.5 and 3.6), and lower bounds in terms of diameter and curvature (Propositions 3.8 and 3.9).
Various graph constructions are discussed in Section 4, namely, the addition of internal edges (Lemma 4.1), combining graphs by bridges (Lemma 4.3), splitting vertices and edges (Lemmas 4.5 and 4.8), taking Cartesian products (Lemma 4.9) and suspensions (Theorem 4.10). In particular, we have $\nu(\tilde{T})=1$ for the suspension $\tilde{T}$ of any tree $T$.
In Section 5 we briefly discuss potential relations between our magneto-spectral height and the spectral gap, which is another widely investigated and fundamental spectral invariant. Examples 5.1 and 5.2 show that $\nu(G)$ cannot be bounded above by the spectral gap (up to any multiplicative scaling). On the positive side, Proposition 5.5 provides evidence that $\nu(G)$ may be bounded below by the spectral gap (up to a multiplicative scaling).
The final Section 6 presents some open problems about the magneto-spectral height which may inspire further work on this topic.
1.2. Context and outlook
It is natural to ask what the magneto-spectral height is actually measuring. This invariant vanishes if and only if the graph does not contain any cycles. Hence this invariant could be viewed as a measure of the complexity of the cycle structure of a graph, where larger cycles seem to contribute less to this complexity than smaller ones.
Magnetic potentials on graphs are special cases of graph connections. A connection on a graph $G=(V,E)$ is a map $\sigma:E^{or}(G)\to O(n)$ or $\sigma:E^{or}(G)\to U(n)$ satisfying $\sigma(y,x)=\sigma(x,y)^{-1}$ for all $(x,y)\in E^{or}(G)$. We refer to $O(n)$ or $U(n)$ as the structure group of the connection. A connection can be viewed as a discrete analogue of a parallel transport in Riemannian Geometry. A connection $\sigma$ gives rise to a connection Laplacian $\Delta^{\sigma}$, defined on function $f:V\to\mathbb{R}^{n}$ or $f:V\to\mathbb{C}^{n}$, with the same definition as in (1). In the case $O(1)=\{-1,1\}$, the associated connection Laplacians agree with the signed Laplacians and in the case $U(1)=S^{1}$, the associated connection Laplacians agree with the magnetic Laplacians. The eigenvalues of connection Laplacians are again all real-valued and non-negative, and we denote the smallest eigenvalue again by $\lambda_{1}^{\sigma}(G)$. Let $\Gamma=O(n)$ or $\Gamma=U(n)$. Our magneto-spectral height has a natural generalization to connection Laplacians via
$\nu_{\Gamma}(G)=\sup_{\sigma}\lambda_{1}^{\sigma}(G),$
where $\sigma$ runs through all connections $\sigma:E^{or}(G)\to\Gamma$. It is interesting to ask whether the $\nu$-invariants of higher dimensional connection Laplacians carry further interesting information about the graph.
$x$$y$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$y_{1}$$y_{2}$$y_{3}$$y_{4}$$y_{5}$$\sigma(x,y)={\xi_{5}}^{2}$
Figure 1. Relation between a directed edge $(x,y)\in E^{or}(G)$ with $\sigma(x,y)={\xi_{5}}^{2}$ and a matching between the fibers $\pi^{-1}(x)$ and $\pi^{-1}(y)$ in the $5$-lift $\widehat{G}$.
Another modification of the $\nu$-invariant can be introduced for cyclic subgroups of $S^{1}$. If we refer to the set of all $k$-th roots of unity as
$S_{k}:=\{{\xi_{k}}^{j}:j\in\{0,1,\dots,k-1\}\}\subset S^{1},\quad\text{with $\xi_{k}=e^{2\pi i/k}$,}$
we can consider the following invariant:
$\nu_{k}(G)=\sup_{\sigma:E^{or}(G)\to S_{k}}\min\left\{\lambda_{1}^{\sigma}(G),\lambda_{1}^{\sigma^{2}}(G),\dots,\lambda_{1}^{\sigma^{k-1}}(G)\right\},$
where $\sigma^{j}(x,y)=(\sigma(x,y))^{j}$ for all oriented edges $(x,y)\in E^{or}(G)$.
In the special case $k=2$, $\nu_{2}(G)$ agrees with the supremum in (2), where the supremum runs only over the signatures instead of all magnetic potentials. Any $\sigma:E^{or}(G)\to S_{k}$ gives rise to a cyclic $k$-lift $\pi:\widehat{G}\to G$ as follows: Every directed edge $(x,y)\in E^{or}(G)$ corresponds to a particular $k$-matching between the fibers $\pi^{-1}(x)=\{x_{1},\dots,x_{k}\}\in\widehat{V}$ and $\pi^{-1}(y)=\{y_{1},\dots,y_{k}\}\in\widehat{V}$ in the lift $\widehat{G}=(\widehat{V},\widehat{E})$, which is determined by $\sigma(x,y)=\xi_{k}^{\ell}\in S_{k}$ as follows $x_{j}\sim_{\widehat{G}}y_{j+\ell}$ for $j\in\{0,1,\dots,k-1\}$, where the indices are taken modulo $k$ (an example of this is illustrated in Figure 1).
The spectrum of the standard Laplacian $\Delta_{\widehat{G}}$ on $\widehat{G}$ is the multiset union of the spectra of the magnetic Laplacians $\Delta^{\sigma^{j}}$ on $G$, for $j=0,1,\dots,k-1$. For further details about cyclic lifts and this spectral relation, we refer the readers to [MSS15], [LPV20] and [ChV17]. These papers are concerned with the question whether bipartite Ramanujan graphs admit cyclic $k$-lifts which are again Ramanujan. In fact, this is true for $k=2,3,4$, as shown in these papers. For $k\geq 5$, bipartite Ramanujan graphs admits also always general $k$-lifts which are again Ramanujan (see [HPS18]), but it is still an open problem, whether this also holds for the smaller class of cyclic $k$-lifts.
1.3. Magneto-spectral heights for Riemannian manifolds
The magneto-spectral height has an analogue in the continuous setting of Riemannian manifolds. While this topic belongs also into the outlook, it is relevant enough to be discussed in a separate subsection. To explain this analogy, we first need to explain some further concepts in the setting of a connected graph $G=(V,E)$: Every magnetic potential $\sigma:E^{or}(G)\to S^{1}$ gives rise to a representation
(3) $\rho_{\sigma}:\pi_{1}(G,x_{0})\to S^{1},$
where $\pi_{1}(G,x_{0})$ is the fundamental group of $G$ with base vertex $x_{0}$. Recall that $\pi_{1}(G,x_{0})$ is the set of all equivalence classes of closed walks in $G$ starting and ending in $x_{0}$. This set has a natural group structure by concatenation and reversion, and the representation (3) is defined as follows:
$\rho_{\sigma}(\gamma)=\sigma(x_{0},x_{1})\sigma(x_{1},x_{2})\cdots\sigma(x_{m-1},x_{m}),$
where $\gamma$ is the equivalence class of the closed walk
$x_{0}\sim x_{1}\sim x_{2}\sim\cdots\sim x_{m-1}\sim x_{m}=x_{0}.$
Moreover, the universal cover of $G$ is a tree $T=(\widetilde{V},\widetilde{E})$ with covering map $\pi:T\to G$, on which the fundamental group $\Gamma=\pi_{1}(G,x_{0})$ acts by deck transformations. Let $\tilde{x}_{0}\in\pi^{-1}(x_{0})\subset\tilde{V}$.
Let $\rho:\Gamma\to S^{1}$ an arbitrary representation. We can identify functions $f\in C(V,\mathbb{C})$ with $\rho$-equivariant functions $\vec{f}\in C(\tilde{V},\mathbb{C})$, that is,
$\vec{f}(\tilde{x}_{0})=f(x_{0})$
and
$\vec{f}(\gamma x)=\rho_{\sigma}(\gamma)\vec{f}(x)\quad\text{for all $x\in\tilde{V}$ and $\gamma\in\Gamma$.}$
The set of all $\rho$-invariant functions on $T$ is denoted by
$C_{\rho}(\widetilde{V},\mathbb{C})=\{g:\widetilde{V}\to\mathbb{C}:g(\gamma x)=\rho(\gamma)g(x)\quad\text{for all $\gamma\in\Gamma$ and $x\in\widetilde{V}$}\},$
and the bijective map $f\mapsto\vec{f}$ is henceforth denoted by $\mathcal{F}_{\rho}:C(V,\mathbb{C})\to C_{\rho}(\widetilde{V},\mathbb{C})$. The we have the following commutative diagram relating the magnetic Laplacian $\Delta^{\sigma}:C(V,\mathbb{C})\to C(V,\mathbb{C})$ and the standard Laplacian $\Delta_{T}:C(\tilde{V},\mathbb{C})\to C(\tilde{V},\mathbb{C})$ on the tree $T$ with $\rho=\rho_{\sigma}$:
[…]
In the special case that the representation $\rho=\rho_{\sigma}$ factors through the finite subgroup $S_{k}$, that is, $\rho:\pi_{1}(G,x_{0})\to S_{k}$, the diagram descends to a corresponding commutative diagram between $\Delta^{\sigma}$ on $G$ and the standard Laplacian $\Delta_{\widehat{G}}$ on the cyclic $k$-lift $\widehat{G}=(\widehat{V},\widehat{E})$ corresponding to $\sigma$:
[…]
where $\mathcal{F}_{\rho}$ maps $f\in C(V,\mathbb{C})$ now to a $\rho$-equivariant function $\vec{f}$ on $\widehat{V}$.
The spectrum of $\Delta_{\widehat{G}}$ is then the multiset disjoint union of the spectra of the magnetic Laplacians $\Delta^{\sigma^{j}}$ on $G$ for $j=0,1,\dots,k-1$.
We now discuss the analogous setting in the case of a compact connected Riemannian manifold $(M,g)$ with base point $x_{0}\in M$. Let $\pi:\widetilde{M}\to M$ be the universal covering map and $\tilde{x}_{0}\in\pi^{-1}(x_{0})\subset\tilde{M}$. Then the fundamental group $\Gamma=\pi_{1}(M,x_{0})$ acts by deck transformations on the universal cover $\widetilde{M}$. Any representation $\rho:\pi_{1}(M,x_{0})\to S^{1}$ gives rise to a flat complex line bundle $L_{\rho}\to M$ and to a (one-dimensional) twisted Laplacian $\Delta^{\rho}:C^{\infty}(M,L_{\rho})\to C^{\infty}(M,L_{\rho})$ with the following properties:
  • (a)
    There exists a one-one correspondence $\mathcal{F}_{\rho}$ between sections $f\in C^{\infty}(M,L_{\rho})$ and $\rho$-equivariant functions $\vec{f}=\mathcal{F}_{\rho}f$ in $C^{\infty}_{\rho}(\widetilde{M},\mathbb{C})$, that is, $\vec{f}$ satisfies
    $\vec{f}(\gamma x)=\rho(\gamma)\vec{f}(x)\quad\text{for all $x\in\widetilde{M}$and $\gamma\in\Gamma$.}$
  • (b)
    We have the following commutative diagram:
[…]
    where $\Delta_{\widetilde{M}}$ is the standard Laplacian on $\widetilde{M}$.
If $\rho$ factors through the finite subgroup $S_{k}$, there exists a finite $k$-cover $\widehat{M}$ of $M$ such that the spectrum of the standard Laplacian $\Delta_{\widehat{M}}$ agrees again with the multiset disjoint union of the spectra of the twisted Laplacians $\Delta^{\rho^{j}}:C^{\infty}(M,L_{\rho^{j}})\to C^{\infty}(M,L_{\rho^{j}})$ on $M$ for $j=0,1,\dots,k-1$. For more information about (higher dimensional) twisted Laplacians see, e.g., [Su89] or [CP23, Chapter 3].
Let us briefly discuss the relation between one-dimensional twisted Laplacians and magnetic Laplacians $\Delta^{\alpha}:C^{\infty}(M,\mathbb{C})\to C^{\infty}(M,\mathbb{C})$ on a compact connected Riemannian manifold $(M,g)$ associated to (real-valued) magnetic potentials $\alpha\in\Omega^{1}(M)$ (see [G24, Definition 1.1] for a real-valued version of twisted Laplacians). The $\alpha$-magnetic Laplacian is given by (see, e.g., [Shi87, (1.1)])
(4) $\Delta^{\alpha}f=\Delta_{M}f-2i\,\langle{\rm{grad}}f,\alpha^{\#}\rangle+\left(i\,d^{*}\alpha+\|\alpha^{\#}\|^{2}\right)f,$
where $\alpha^{\#}$ is the vector field determined by the relation $\alpha(Y)=\langle Y,\alpha^{\#}\rangle$ and $d^{*}:\Omega^{1}(M)\to C^{\infty}(M)$ is the formal adjoint of the exterior derivative $d:C^{\infty}(M)\to\Omega^{1}(M)$. The operator $\Delta^{\alpha}$ on a compact Riemannian manifold is self-adjoint with non-negative discrete spectrum, and it can also be viewed as the operator $\Delta^{\tilde{\alpha}}$, acting on $\Gamma$-periodic functions $\tilde{f}\in C_{\Gamma}(\tilde{M},\mathbb{C})$ with $\tilde{\alpha}=\pi^{*}\alpha\in\Omega^{1}(\widetilde{M})$.
Any harmonic magnetic potential $\alpha\in\Omega^{1}(M)$ (that is, $d\alpha=d^{*}\alpha=0$) gives rise to a representation $\rho_{\alpha}:\pi_{1}(M,x_{0})\to S^{1}$ as follows:
$\rho_{\alpha}(\gamma)=\exp(-i\,\int_{\tilde{x}_{0}}^{\gamma\tilde{x}_{0}}\tilde{\alpha}),$
where the path integral $\int_{\tilde{x}_{0}}^{\gamma\tilde{x}_{0}}\tilde{\alpha}$ is independent on the path by Stokes and the fact that $d\tilde{\alpha}=0$. Moreover, for any harmonic magnetic potential $\alpha$ and associated representation $\rho=\rho_{\alpha}$, the operators $\Delta^{\tilde{\alpha}}:C_{\Gamma}^{\infty}(\widetilde{M},\mathbb{C})\to C_{\Gamma}^{\infty}(\widetilde{M},\mathbb{C})$ and $\Delta^{\rho}:C_{\rho}^{\infty}(\widetilde{M},\mathbb{C})\to C_{\rho}^{\infty}(\widetilde{M},\mathbb{C})$ are unitarily equivalent via the gauge function $\tau\in C^{\infty}(\widetilde{M},S^{1})$, given by
$\tau(\tilde{x})=\exp\left(-i\int_{\tilde{x}_{0}}^{\tilde{x}}\tilde{\alpha}\right).$
This follows from ${\rm{grad}}\,\tau=-i\tilde{\alpha}^{\#}$, $\Delta_{\widetilde{M}}\,\tau=\|\tilde{\alpha}^{\#}\|^{2}$, $d^{*}\tilde{\alpha}=0$ and
$\tau^{-1}\Delta_{\widetilde{M}}(\tau\tilde{f})=\Delta_{\widetilde{M}}\tilde{f}-2i\langle{\rm{grad}}\tilde{f},\tilde{\alpha}^{\#}\rangle+\|\tilde{\alpha}^{\#}\|^{2}\tilde{f}=\Delta^{\tilde{\alpha}}\tilde{f},$
for any $\Gamma$-periodic function $\tilde{f}$. This shows $\tau^{-1}\Delta^{\rho}\tau=\Delta^{\tilde{\alpha}}$ and, therefore, every magnetic Laplacian with a harmonic potential is a one-dimensional twisted Laplacian. However, the converse is not true: There are no harmonic $1$-forms on the real projective plane $\mathbb{R}P^{2}$, but there is a non-constant representation $\rho:\pi_{1}(\mathbb{R}P^{2},x_{0})\to\{\pm 1\}$.
Our magneto-spectral invariant in the manifold setting is defined as follows:
Definition 1.3 (Magneto-spectral height for manifolds).
Let $(M,g)$ be a compact Riemannian manifolds. The magneto-spectral height of $M$ is
$\nu(M)=\sup_{\rho}\lambda_{1}^{\rho}(M),$
where $\rho:\pi_{1}(M,x_{0})\to S^{1}$ runs through all $S^{1}$-representations and $\lambda_{1}^{\rho}(M)$ is the smallest eigenvalue of the twisted Laplacian $\Delta^{\rho}$.
It follows directly from its definition, that we have $\nu(M)=0$ for all simply connected manifolds $M$. This is the analogue of the property $\nu(T)=0$ for finite trees. Let us finish this subsection with the following two questions.
Question 1.4.
Another potential continuous counterpart of the magneto-spectral height for compact Riemannian manifolds $(M,g)$ could be
(5) $\sup_{\alpha\in\Omega^{1}(M)}\lambda_{1}^{\alpha}(M),$
where $\alpha$ runs through all magnetic potentials on $M$ and $\lambda_{1}^{\alpha}(M)$ is the smallest eigenvalue of the magnetic Laplacian $\Delta^{\alpha}$, given in (4). Readers may ask why we did not choose this alternative definition.
Since, by Hodge Theory, every $\alpha\in\Omega^{1}(M)$ can be uniquely written as the sum $\alpha_{0}+df+d^{*}\omega$ with harmonic $\alpha_{0}\in\Omega^{1}(M)$, $f\in C^{\infty}(M)$ and $\omega\in\Omega^{2}(M)$, the term $df$ can be gauged away and (5) agrees therefore with
$\sup_{\alpha_{1}}\lambda_{1}^{\alpha_{1}}(M),$
where $\alpha_{1}$ runs through all magnetic potentials on $M$ with $d^{*}\alpha_{1}=0$. However, any term $d^{*}\omega\neq 0$ in the Hodge decomposition of $\alpha$ cannot be gauged away, which implies that we cannot restrict the supremum to run only over harmonic magnetic potentials.
Due to this fact, this alternative definition has the following significant disadvantage: It is no longer true that the expression (5) vanishes for simply connected compact Riemannian manifolds. Even worse, the expression (5) is infinite for the simply connected $3$-sphere (see [CGHP25, Section 5.2]).
Question 1.5.
In the continuous setting of manifolds, Ramanujan graphs correspond to hyperbolic surfaces $S$ without small eigenvalues, that is $\lambda_{1}(S)\geq 1/4$. It is natural to ask, whether the lower Ramanujan bound $\nu(G)\geq d-2\sqrt{d-1}$ for $d$-regular graphs has an continuous analogue, that is, whether $\nu(S)\geq 1/4$ for all compact (or finite volume) hyperbolic surfaces. A closely related question is whether all compact (or finite volume) hyperbolic surfaces admit finite covers with no new eigenvalues in the interval $[0,1/4)$. For significant recent progress in this question, see [HM23] and the references therein.
2. Examples of graphs
We start this section about examples with the class of graphs whose magneto-spectral height vanishes.
Lemma 2.1.
Any magnetic potential can be gauged away if and only if $G=(V,E)$ is a forest. Therefore, we have $\nu(G)=0$ if and only if $G$ is a forest.
2.1. Cycles
Example 2.2 (The cycle $C_{d}$).
Let $d\geq 3$. We will show for the $d$-cycle $C_{d}$ that we have
(6) $\nu(C_{d})=2-2\cos(\pi/d).$
We enumerate the vertices of $C_{d}$ by $x_{0},x_{1},\dots,x_{n-1}$ and impose the non-triviality of $\sigma$ only in the edge $\{x_{n-2},x_{n-1}\}$, that is
$\sigma_{t}(x_{n-2},x_{n-1})=e^{it}.$
The Laplacian $\Delta^{\sigma_{t}}$ is then represented by the following matrix:
$\Delta^{\sigma_{t}}\cong\begin{pmatrix}\begin{array}[]{ccccccc|c}2&-1&0&\cdots&\cdots&0&0&-1\\ -1&2&-1&&&0&0&0\\ 0&-1&2&\ddots&&0&0&0\\ \vdots&&\ddots&\ddots&\ddots&&&\vdots\\ \vdots&&&\ddots&\ddots&\ddots&&\vdots\\ 0&0&0&&\ddots&2&-1&0\\ 0&0&0&&&-1&2&-e^{it}\\ \hline\cr-1&0&0&\cdots&\cdots&0&-e^{-it}&2\end{array}\end{pmatrix},$
and the eigenfunctions $f_{j}:C_{d}\to\mathbb{C}$ are illustrated in Figure 2 with $\xi_{d}=e^{2\pi i/d}$. Note that, in the case $t=0$, the red factors disappear and we end up with the eigenfunctions of the standard Laplacian on $C_{d}$. Moreover, we check for any $j_{1}\neq j_{2}$ that
$\langle f_{j_{1}},f_{j_{2}}\rangle$ $=\sum_{\ell=0}^{d-2}\left(e^{\frac{it}{d}}\xi_{d}^{j_{1}}\right)^{\ell}\overline{\left(e^{\frac{it}{d}}\xi^{j_{2}}_{d}\right)^{\ell}}+\left(e^{\frac{it}{d}}\xi_{d}^{j_{1}}\right)^{-1}\overline{\left(e^{\frac{it}{d}}\xi_{d}^{j_{2}}\right)^{-1}}$
(7) $=\sum_{\ell=-1}^{d-2}\xi_{d}^{(j_{1}-j_{2})\ell}=\xi_{d}^{-(j_{1}-j_{2})}\sum_{\ell=0}^{d-1}\xi_{d}^{(j_{1}-j_{2})\ell}=0.$
It is easy to check that the corresponding eigenvalue to $f_{j}$ is
(8) $\mu_{j}(t)=4\sin^{2}\left(\frac{t+2\pi j}{2d}\right)=2-2\cos\left(\frac{t+2\pi j}{d}\right)=2-\left(\xi_{d}^{j}e^{i\frac{t}{d}}+\xi_{d}^{-j}e^{-i\frac{t}{d}}\right).$
Note that we have for $t\in[0,\pi]$,
$\cos\left(\frac{t}{d}\right)\geq\cos\left(\frac{\pi}{d}\right)\geq\cos\left(\frac{t+2j\pi}{d}\right),$
since $\frac{t+2\pi j}{d}\in[\frac{2\pi}{d},2\pi-\frac{\pi}{d}]$ for $j=1,2,\dots,d-1$. This implies for $t\in[0,\pi]$ that
$\mu_{0}(t)\leq\mu_{j}(t)\quad\text{for all $j=1,2,\dots,d-1$},$
with equality if and only if $t=\pi$ and $j=d-1$. Moreover, we have for $t\in[0,\pi]$,
$\lambda_{1}(\Delta^{\sigma_{t}})=\mu_{0}(t)=2-2\cos(t/d).$
Since the eigenvalue graphs are symmetric with respect to $\pi$, we conclude that
$\nu(C_{d})=\mu_{0}(\pi)=2-2\cos(\pi/d).$
Moreover, the multiplicity of the smallest eigenvalue of $\Delta^{\sigma_{\pi}}$ is precisely $2$, since for $t=\pi$,
$\mu_{0}(\pi)=\mu_{d-1}(\pi)<\mu_{j}(\pi)\quad\text{for all $j=1,2,\dots,d-2$.}$
[…]
Figure 2. Eigenfunctions $f_{j}$ of the magnetic Laplacian $\Delta^{\sigma}$ on $C_{d}$ with $j=0,1,\dots,d-1$. The red factors describe the influence of the magnetic potential $\sigma_{t}$.
The previous example directly implies the following asymptotics of $\nu(C_{d})$ as $d\to\infty$.
Corollary 2.3.
We have
$\nu(C_{d})=4\sin^{2}(\pi/(2d))\leq\frac{\pi^{2}}{d^{2}},$
and therefore $\nu(C_{d})\to 0$ as $d\to\infty$.
2.2. Complete graphs
The following lemma can be viewed as a special case of the later Theorem 3.3(a) in Section 3. For the reader’s convenience, we also present a proof in the simpler case of this lemma.
Lemma 2.4.
Let $G=(V,E)$ be a graph and $T\subset V$ be a subgraph which is a forest. Then we have
(9) $\nu(G)\leq\frac{|\partial T|}{|T|},$
where $\partial T\subset E$ is the set of edges connecting vertices of $T$ with vertices of $V\setminus T$. In particular, we have
(10) $\nu(G)\leq\min_{x\sim y}\frac{d_{x}+d_{y}}{2}-1.$
Remark 2.5.
Inequality (9) in Lemma 2.4 can also be viewed as the inequality
$\nu(G)\leq h_{\rm{forest}}(G),$
where $h_{\rm{forest}}$ is a Cheeger-type constant, which takes the infimum of isoperimetric ratios only for subgraphs which are forests.
Proof.
We first present the proof in the special case when $T$ is a tree. Let $\sigma:E^{or}\to S^{1}$ be an arbitrary magnetic potential and $x_{0}\in T$. Choose $f:V\to\mathbb{C}$ as follows: We set $f(x_{0})=1$ For any vertex $x\in T\setminus\{x_{0}\}$, choose the unique path $x_{0}\sim x_{1}\sim\cdot\sim x_{l}=x$ in $T$ and define
$f(x)=\sigma(x_{0},x_{1})^{-1}\sigma(x_{1},x_{2})^{-1}\cdots\sigma(x_{l-1},x_{l})^{-1}.$
For any vertex $x\in V\setminus T$, define
$f(x)=0.$
Then we have
$\lambda_{1}^{\sigma}(G)\leq\mathcal{R}^{\sigma}(f)=\frac{\displaystyle{\sum_{x\sim y\atop x\in T,y\in V\setminus T}}|f(x)-\sigma(x,y)f(y)|^{2}}{\sum_{x\in T}|f(x)|^{2}}=\frac{|\partial T|}{|T|}.$
In the case that $T$ is a forest, we choose in every connected component of $T$ a base point $x_{0}$ to define the test function $f$ and argue accordingly.
For the second part of the proposition, choose $T=\{x,y\}$ for any pair $x\sim y$. ∎
Example 2.6 (The complete graph $K_{d}$).
For the complete graph $K_{d}$, we will show that
$\nu(K_{d})=d-2.$
We have
$\lambda^{\overline{\sigma}}(K_{d})=d-2$
for the anti-balanced magnetic potential $\overline{\sigma}\equiv-1$, since
$\Delta^{\overline{\sigma}}\equiv(d-2){\rm{Id}}_{d}+J_{d},$
with the all-$1$ square matrix $J_{d}$ of size $d$, and all non-zero vectors perpendicular to the all-$1$ vector of size $d$ are eigenvectors to the eigenvalue $d-2$, whereas the all-$1$ vector of size $d$ is an eigenvector to the eigenvalue $22(d-1)>d-$ 2. Since $K_{d}$ is $(d-1)$-regular, it follows from Lemma 2.4 that
$\nu(K_{d})\leq\min_{x\sim y}\frac{d_{x}+d_{y}}{2}-1=(d-1)-1=d-2.$
Since $\lambda^{\overline{\sigma}}(K_{d})=d-2$, we have equality $\nu(G)=d-2$, assumed by the anti-balanced magnetic potential $\overline{\sigma}$. For this potential, the multiplicity of $\lambda_{1}^{\overline{\sigma}}$ has multiplicity $d-2$.
To derive particular rigidity results for complete graphs (Corollary 2.8 and Proposition 2.9 below), we start with the following general fact.
Proposition 2.7.
Let $G=(V,E)$ be a graph and $\{x,y\}\in E$ be an edge. If
$\nu(G)=\frac{d_{x}+d_{y}}{2}-1,$
then we have
$S_{1}(x)\setminus\{y\}=S_{1}(y)\setminus\{x\}.$
In particular, we have $d_{x}=d_{y}$.
Proof.
Let $\hat{\sigma}:E^{or}(G)\to S^{1}$ be a maximal magnetic potential, that is,
$\nu(G)=\lambda_{1}^{\hat{\sigma}}(G).$
Let $f:V\to\mathbb{C}$ be the following function
$f(z)=\begin{cases}1&\text{if $z=x$,}\\ \hat{\sigma}(x,y)^{-1}&\text{if $z=y$,}\\ 0&\text{otherwise.}\end{cases}$
Then we have
$\frac{d_{x}+d_{y}}{2}-1=\nu(G)=\lambda_{1}^{\hat{\sigma}}(G)\leq\mathcal{R}^{\hat{\sigma}}(f)=\frac{d_{x}+d_{y}}{2}-1.$
Therefore, $f$ is an eigenfunction of $\Delta^{\hat{\sigma}}$ to the eigenvalue $\frac{d_{x}+d_{y}}{2}-1$. Let us denote the neighbours of $x$ by $y=y_{1},y_{2},\dots,y_{d_{x}}$. Assume $y_{j}$, $j\in\{2,\dots,d_{x}\}$ is not a neighbour of $y=y_{1}$. Then we have
$\Delta^{\hat{\sigma}}f(y_{j})=\sum_{z\sim y_{j}}f(y_{j})-\hat{\sigma}(y_{j},z)f(z)=-\hat{\sigma}(y_{j},x)f(x)=0,$
which is a contradiction. This shows that all neighbours of $x$, different from $y$, are also neighbours of $y$. Interchanging the role of $x$ and $y$ implies that we have $d_{x}=d_{y}.$
Now we can prove the following rigidity result.
Corollary 2.8.
Any connected graph $G=(V,E)$ satisfying
$\nu(G)=\max_{x\sim y}\frac{d_{x}+d_{y}}{2}-1$
is already a complete graph. In particular, any connected graph $G$ with
$\nu(G)=d_{\max}-1,$
where $d_{\max}$ is the maximal vertex degree of $G$, is a complete graph $K_{d_{\max}+1}$.
Proof.
It follows from Lemma 2.4 that we have
$\min_{x\sim y}\frac{d_{x}+d_{y}}{2}-1\geq\nu(G)=\max_{x\sim y}\frac{d_{x}+d_{y}}{2}-1,$
which implies $d_{x}+d_{y}$ is constant for any edge $\{x,y\}\in E$. It follows from Proposition 2.7 that any connected component of $G$ is regular and complete. Since $G$ is assumed to be connected, the first statement of the corollary follows.
Since
$\max_{x\sim y}\frac{d_{x}+d_{y}}{2}\leq d_{\max},$
the condition $\nu(G)=d_{\max}-1$ implies that we have $d_{x}=d_{y}=d_{\max}$ for all adjacent vertices. Since $G$ is connected, $G$ is regular and Proposition 2.7 implies again that $G$ is a complete graph. ∎
We also have the following rigidity result.
Proposition 2.9.
Let $\hat{\sigma}:E^{or}(K_{d})\to S^{1}$ be a maximal magnetic potential of the complete graph $K_{d}$. Then $\hat{\sigma}$ is automatically gauge equivalent to the anti-balanced signature $\overline{\sigma}\equiv-1$.
Proof.
Let $\hat{\sigma}$ be a maximal magnetic potential on $K_{d}$. For $k,l\in\{1,\dots,d\}$, we introduce the function
$f_{k,l}(x_{m})=\begin{cases}1&\text{if $m=k$,}\\ \hat{\sigma}(x_{k},x_{l})^{-1}&\text{if $m=l$,}\\ 0&\text{otherwise.}\end{cases}$
Then we have
$\lambda_{1}^{\hat{\sigma}}(K_{d})=d-2=\mathcal{R}^{\hat{\sigma}}(f_{k,l}),$
and $f_{k,l}$ is therefore an eigenfunction of $\Delta^{\hat{\sigma}}$ to the eigenvalue $d-1$. The eigenfunction equation yields for $m\not\in\{k,l\}$,
$\Delta^{\hat{\sigma}}f_{k,l}(x_{m})=\underbrace{\sum_{j\neq m}f_{k,l}(x_{m})-\hat{\sigma}(x_{m},x_{j})f_{k,l}(x_{j})}_{\hat{\sigma}(x_{m},x_{k})-\hat{\sigma}(x_{m},x_{l})\hat{\sigma}(x_{k},x_{l})^{-1}}=(d-1)f_{k,l}(x_{m})=0,$
which implies
$\hat{\sigma}(x_{m},x_{k})\hat{\sigma}(x_{k},x_{l})\hat{\sigma}(x_{l},x_{m})=-1.$
Choosing the gauge function $\tau:V\to S^{1}$,
$\tau(z)=\begin{cases}-1,&\text{if $z=x_{1}$,}\\ \hat{\sigma}(x_{1},x_{j})^{-1},&\text{if $z=x_{j}$, $j\geq 2$,}\end{cases}$
we obtain
$\hat{\sigma}^{\tau}(x_{1},x_{j})=\tau(x_{1})^{-1}\hat{\sigma}(x_{i},x_{j})\tau(x_{j})=-1\quad\text{for all $j\geq 2$.}$
Since, for $k,l\geq 2$, $k\neq l$,
$\hat{\sigma}^{\tau}(x_{k},x_{l})=\hat{\sigma}^{\tau}(x_{1},x_{k})\hat{\sigma}^{\tau}(x_{k},x_{l})\hat{\sigma}^{\tau}(x_{l},x_{1})=\hat{\sigma}(x_{1},x_{k})\hat{\sigma}(x_{k},x_{l})\hat{\sigma}(x_{l},x_{1})=-1,$
$\hat{\sigma}^{\tau}$ is the anti-balanced signature. ∎
2.3. Wheel graphs
Example 2.10 (Suspension of a tree).
Let $T$ be a tree and $\tilde{T}$ be its suspension, that is, we add a vertex $x$ to $T$ and connect it to all vertices in $T$. Then we have
$\nu(\tilde{T})\leq 1.$
This follows directly from (9) in Lemma 2.4. In fact, we will see later (see Theorem 4.10) that we have $\nu(\tilde{T})=1$ for all trees.
Next we consider suspensions of cycles. The suspension of a cycle $C_{d}$ ($d\geq 3$) is the wheel graph $W_{d}$. In the next example we compute an explicit lower bound for $\nu(W_{d})$. In contrast to the suspension of trees, we will see that $\nu(W_{d})>1$.
Example 2.11 (The wheel graphs $W_{d}$).
Let $d\geq 3$ and $x_{0},\dots,x_{d-1}$ be the vertices of $W_{d}$ in the cycle $C_{d}\subset W_{d}$ and $x\in W_{d}$ be the suspension vertex. Let $\sigma_{1}:E^{or}(W_{d})\to\{-1,1\}$ be given by
$\sigma_{1}(x_{j},x_{j+1})=\begin{cases}1&\text{for $j=0,1,\dots,d-2$,}\\ -1&\text{for (j,j+1)=(d-1,0),}\end{cases}$
and
$\sigma_{1}(x,x_{j})=\begin{cases}-\left(e^{-i\frac{2k+1}{d}\pi}\right)^{j}&\text{for $j=0,1,\dots,d-2$,}\\ -e^{i\frac{2k+1}{d}\pi}&\text{for $j=d-1$.}\end{cases}$
We will show that the eigenvalues of $\Delta^{\sigma_{1}}$ on $W_{d}$ (with multiplicities) are given by the values
$3-2\cos\left(\frac{\pi+2\pi j}{d}\right)\quad\text{with $j=0,1,\dots,d-1$, $j\neq k$},$
all counted once, and the roots of
$\lambda^{2}-\left(d+3-2\cos\left(\frac{2k+1}{d}\pi\right)\right)\lambda+\left(2-2\cos\left(\frac{2k+1}{d}\pi\right)\right)d=0,$
(with $d=2k$ for even $d$ and $d=2k+1$ for odd $d$) which are also counted once. Note that this equation simplifies in the case $d=2k+1$ to
$\lambda^{2}-(d+5)\lambda+4d=0.$
Using Lemma 2.12 below, this implies that
$\nu(W_{d})\geq\lambda_{1}^{\sigma_{1}}(W_{d})=3-2\cos\left(\frac{\pi}{d}\right)=1+\nu(C_{d}).$
Note that the restriction of $\sigma_{1}$ to $C_{d}$, denoted again by $\sigma_{1}$, for simplicity, is a maximal magnetic potential of $C_{d}$. The eigenfunctions of $\Delta^{\sigma_{1}}_{C_{d}}$ are given by (cf. Figure 2 with $t=\pi$)
$f_{j}=\begin{pmatrix}1\\ e^{\frac{i\pi}{d}}\xi_{d}^{j}\\ (e^{\frac{i\pi}{d}}\xi_{d}^{j})^{2}\\ \vdots\\ (e^{\frac{i\pi}{d}}\xi_{d}^{j})^{d-2}\\ e^{-\frac{i\pi}{d}}(\xi_{d}^{j})^{-1}\end{pmatrix},\quad\text{ $j=0,1,\cdots,d-1$.}$
Note that we have
$\overline{f_{k}}=\begin{pmatrix}-\sigma_{1}(x,x_{0})\\ -\sigma_{1}(x,x_{1})\\ -\sigma_{1}(x,x_{2})\\ \vdots\\ -\sigma_{1}(x,x_{d-2})\\ -\sigma_{1}(x,x_{d-1})\end{pmatrix},$
and, therefore, we have
$\Delta^{\sigma_{1}}_{W_{d}}\cong\begin{pmatrix}\begin{array}[]{c|ccc}d&&\overline{f_{k}}^{\top}&\\ \hline\cr&&&\\ f_{k}&&\Delta^{\sigma_{1}}_{C_{d}}+\mathrm{Id}&\\ &&&\end{array}\end{pmatrix}.$
The orthogonality relation (2.2) implies that $\Delta^{\sigma_{1}}_{W_{d}}$ has the following set of $(d-1)$ eigenfunctions:
$\left\{\begin{pmatrix}0\\ f_{j}\end{pmatrix},\,\,\,j=0,1,\cdots,k-1,k+1,\cdots,d-1\right\},$
corresponding to the eigenvalues $\mu_{j}(\pi)+1$, $j\neq k$, with the eigenvalues $\mu_{j}(\pi)$ of $\Delta_{C_{d}}^{\sigma_{1}}$ given in (8). The remaining $2$-dimensional vector space orthogonal to $span\{f_{0},f_{1},\cdots,f_{k-1},f_{k+1},\cdots,f_{d-1}\}$ can be written as
$\left\{\begin{pmatrix}a\\ bf_{k}\end{pmatrix},\,\,\,a,b\in\mathbb{C}\right\}.$
Solving the eigenfunction equation
$\begin{pmatrix}d&\overline{f_{k}}^{\top}\\ f_{k}&\Delta^{\hat{\sigma}}_{C_{d}}+\mathrm{Id}\end{pmatrix}\begin{pmatrix}a\\ bf_{k}\end{pmatrix}=\lambda\begin{pmatrix}a\\ bf_{k}\end{pmatrix}$
leads to
$\begin{cases}(\lambda-d)a=db,\quad&\text{(A)}\\ a=(\lambda-\mu_{k}(\pi)-1)b,&\text{(B)}\end{cases}$
with
$\mu_{k}(\pi)=2-2\cos\left(\frac{2k+1}{d}\pi\right).$
Replacing $a$ with $(B)$ in $(A)$ yields
$(\lambda-d)(\lambda-\mu_{k}(\pi)-1)b=db.$
Since $b=0$ implies $\begin{pmatrix}a\\ bf_{k}\end{pmatrix}=0$, in contradiction to an eigenfunction, we can divide by $b\neq 0$ and obtain the quadratic equation
$\lambda^{2}-(d+\mu_{k}(\pi)+1)\lambda+\mu_{k}(\pi)d=0,$
which gives the remaining two eigenvalues of $\Delta^{\sigma_{1}}_{W_{d}}$.
Lemma 2.12.
  • (a)
    Let $d\geq 3$ be odd. Then we have
    $\frac{d+5-\sqrt{(d+5)^{2}-16d}}{2}\geq 3-2\cos(\pi/d).$
  • (b)
    Let $d\geq 4$ be even. Then we have
    $\frac{d+3+2\cos(\pi/d)-\sqrt{(d+3+2\cos(\pi/d))^{2}-4d(2+2\cos(\pi/d))}}{2}\geq 3-2\cos(\pi/d).$
Proof.
The inequality in (a) is an equality in the case $d=3$. It is also easy to verify (b) for $d=4$. So we can restrict to the case $d\geq 5$.
Both (a) and (b) are direct consequences of the following two facts:
  • (i)
    We have
    $d+3+2c-\sqrt{(d+3+2c)^{2}-4d(2+2c)}\geq 6-4c$
    for $d\geq 5$ and $c=\cos(\pi/d)$.
  • (ii)
    The map
    $[0,1]\ni c\mapsto d+3+2c-\sqrt{(d+3+2c)^{2}-4d(2+2c)}$
    is monotone increasing for all $d\geq 5$.
(i) is equivalent to
$(4c-1)d+8c^{2}-12c\geq 0$
for $d\geq 5$ and $c=\cos(\pi/d)$. Since $d\geq 5$, we have
$c=\cos(\pi/d)\geq\cos(\pi/4)=\frac{\sqrt{2}}{2}.$
So it suffices to show
$(4c-1)d+8c^{2}-12c\geq 0$
for $d\geq 5$ and $c\in[\sqrt{2}/2,1]$. Since $04c-1>$ 0, the left hand side is increasing in $d$, and it suffices to show the inequality for $d=5$. This reduces the problem to show that $58c^{2}+8c-$ 5 is non-negative for $c\in[\sqrt{2}/2,1]$, which is obviously true. This finishes the proof of (i).
To prove (ii), we set $\alpha=3+2c$ and need to verify, for all $d\geq 5$, that
$\alpha\mapsto d+\alpha-\sqrt{(d+\alpha)^{2}-4d(\alpha-1)}=d+\alpha-\sqrt{(d-\alpha)^{2}+4d}$
is monotone increasing for $\alpha\in[3,5]$. This is obviously the case, which completes the proof of the lemma. ∎
The following result states that our lower bound for $\nu(W_{d})$ is sharp.
Proposition 2.13.
For any wheel graph $W_{d}$, $d\geq 3$, we have
(11) $\nu(W_{d})=\nu(C_{d})+1=3-2\cos(\pi/d).$
Proof.
We apply Theorem 3.3 in Section 3 below to wheel graphs, by choosing $G=(V,E)$ as $W_{d}$ and the subgraph $G_{0}=(V_{0},E_{0})$ as the cycle $C_{d}$. Then we obtain the inequality:
$\nu(W_{d})\leq\nu(C_{d})+1.$
Combining this with the results in Example 2.11 and Example 2.2 completes the proof of the proposition. ∎
2.4. Cospectral graphs
It is known (see [ZLY09]) that wheel graphs $W_{k}$, $k\neq 6$, are uniquely determined by their Laplace spectrum. However, there exists a cospectral graph $\widehat{G}$ for the wheel graph $W_{6}$. Both graphs are illustrated in Figure 3.
[…]
Figure 3. The two isospectral graphs $W_{6}$ and $\widehat{G}$
The matrix representations of the standard Laplacian on $W_{6}$ is of the form
$\Delta_{W_{6}}\cong\begin{pmatrix}6&-1&-1&-1&-1&-1&-1\\ -1&3&-1&0&0&0&-1\\ -1&-1&3&-1&0&0&0\\ -1&0&-1&3&-1&0&0\\ -1&0&0&-1&3&-1&0\\ -1&0&0&0&-1&3&-1\\ -1&-1&0&0&0&-1&3\end{pmatrix},$
and its spectrum is given by
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$ $\lambda_{6}$ $\lambda_{7}$
00 $2$ $2$ $4$ $4$ $5$ $7$
The standard Laplacian $\Delta_{\widehat{G}}$ on $\widehat{G}$ has the same spectrum, and the matrix representation of the signed Laplacian $\Delta_{\widehat{G}}^{\sigma}$ with signature $\sigma:E^{or}(\widehat{G})\to\{-1,1\}$ assuming the value $+1$ on the red edges and the value $-1$ on the blue edges in Figure 3 is of the form
$\Delta_{\widehat{G}}^{\sigma}\cong\begin{pmatrix}5&0&-1&-1&-1&1&1\\ 0&5&1&1&-1&-1&-1\\ -1&1&3&1&0&0&0\\ -1&1&1&3&0&0&0\\ -1&-1&0&0&2&0&0\\ 1&-1&0&0&0&3&1\\ 1&-1&0&0&0&1&3\end{pmatrix}.$
The spectrum of $\Delta_{\widehat{G}}^{\sigma}$ is given by
$\lambda_{1}^{\sigma}$ $\lambda_{2}^{\sigma}$ $\lambda_{3}^{\sigma}$ $\lambda_{4}^{\sigma}$ $\lambda_{5}^{\sigma}$ $\lambda_{6}^{\sigma}$ $\lambda_{7}^{\sigma}$
$\frac{7}{2}-\frac{\sqrt{17}}{2}$ $\frac{9}{2}-\frac{\sqrt{33}}{2}$ $2$ $2$ $4$ $\frac{7}{2}+\frac{\sqrt{17}}{2}$ $\frac{9}{2}+\frac{\sqrt{33}}{2}$
Note that we have
$\nu(W_{6})=3-2\cos(\pi/6)\approx 1.2679<1.4384\approx\frac{7}{2}-\frac{\sqrt{17}}{2}\leq\nu(\widehat{G}).$
This shows that the cospectral graphs $W_{6}$ and $\widehat{G}$ can be distiguished by the magneto-spectral height.
2.5. Some Counterexamples
Question 2.14.
The $\lambda_{1}$-muliplicity considerations in Examples 2.2 and 2.6 motivates the question whether the $\lambda_{1}^{\hat{\sigma}}(G)$-eigenspace of a maximal magnetic potential $\hat{\sigma}$ has always dimension $\geq 2$ if $\nu(G)>0$. Example 2.14 shows that this is wrong.
Example 2.15 (Counterexample to Question 2.14).
[…]
Figure 4. A counterexample to Question 2.14
We consider a triangle with a dangling edge added, as shown in Figure 4. The magnetic potential is determined by one parameter $t\in[0,2\pi]$ and denoted as $\sigma_{t}$, and the Laplacian matrix is of the form
$\Delta^{\sigma_{t}}\cong\begin{pmatrix}1&-1&0&0\\ -1&3&-1&-1\\ 0&-1&2&-e^{it}\\ 0&-1&-e^{-it}&2\end{pmatrix}.$
Hence, the characteristic polynomial is given by
$(\lambda-1)(\lambda^{3}-7\lambda^{2}+12\lambda-2(1-\cos(t))).$
The maximal smallest root appears at the choice $t=\pi$, see the eigenvalues picture in Figure 5, in which case we have the characteristic polynomial as follows:
$(\lambda-1)(\lambda^{3}-7\lambda^{2}+12\lambda-4)=(\lambda-1)(\lambda-2)(\lambda^{2}-5\lambda+2),$
and $\nu(G)=\frac{5-\sqrt{17}}{2}\approx 0.438447187$. In addition, this figure shows that a maximal magnetic potential of a triangle with a dangling edge added is the anti-balanced signature $\overline{\sigma}\equiv-1$ and the smallest eigenvalue $\lambda^{\overline{\sigma}}_{1}(G)$ has multiplicity $1$.
Figure 5. The functions $t\mapsto\lambda_{1}^{\sigma_{t}}(G),\lambda_{2}^{\sigma_{t}}(G),\lambda_{3}^{\sigma_{t}}(G),\lambda^{\sigma_{t}}_{4}(G)$
Question 2.16.
Note that, in the examples $G=C_{d}$ and $G=K_{d}$, there exists a magnetic potential $\hat{\sigma}$ only assuming values ±$\pm 1$, which satisfies
$\nu(G)=\lambda_{1}^{\hat{\sigma}}(G).$
Therefore, it is natural to ask whether, for any graph $G=(V,E)$, there exists a magnetic potential $\hat{\sigma}:E^{or}(G)\to\{-1,1\}$ satisfying
$\lambda_{1}^{\hat{\sigma}}(G)=\nu(G).$
Example 2.17 shows that this is not always the case.
Example 2.17 (Counterexample to Question 2.16).
We consider a graph and a corresponding magnetic potential as illustrated in Figure 6. Note that this graph is the suspension of a path of length $2$. The magnetic potential is determined by two parameters $\alpha,\beta\in[0,2\pi]$ and denoted by $\sigma_{\alpha,\beta}$, and the Laplacian is given by the matrix
$\Delta^{\sigma_{\alpha,\beta}}\cong\begin{pmatrix}3&-1&-1&-1\\ -1&2&0&-e^{i\alpha}\\ -1&0&2&-e^{i\beta}\\ -1&-e^{-i\alpha}&-e^{-i\beta}&3\end{pmatrix}.$
Its characteristic polynomial is given by
$t^{4}-10t^{3}+32t^{2}-(36-2(\cos\alpha+\cos\beta))t+10-4(\cos\alpha+\cos\beta)-2\cos(\alpha-\beta).$
[…]
Figure 6. A counterexample to Question 2.16
The function $(\alpha,\beta)\mapsto\lambda_{1}^{\sigma_{\alpha,\beta}}(G)$ is illustrated in Figure 7. For example, we have the following eigenvalues for all magnetic potentials only assuming values $\{-1,1\}$:
$\lambda_{1}^{\sigma_{0,0}}(G)$ $=0,$
$\lambda_{1}^{\sigma_{0,\pi}}(G)$ $=\lambda_{1}^{\sigma_{\pi,0}}(G)=0.5857864376,$
$\lambda_{1}^{\sigma_{\pi,\pi}}(G)$ $=0.7639320225.$
However, we have
$\lambda_{1}^{\sigma_{2\pi/3,4\pi/3}}(G)=\lambda_{1}^{\sigma_{4\pi/3,2\pi/3}}(G)=1.$
The magnetic potentials $\sigma$ with values $\{-1,1\}$ provide therefore clearly not the maximal first magnetic eigenvalue $\nu(G)$.
Figure 7. The function $(\alpha,\beta)\mapsto\lambda_{1}^{\sigma_{\alpha,\beta}}(G)$ 2.16
Question 2.18.
Proposition 3.1 below states that $\nu(G)$ is bounded above by the average vertex degree
$\frac{1}{|V|}\sum_{x\in V}d_{x}=\frac{2|E|}{|V|}.$
It is therefore natural to ask, whether $\nu(G)$ is also bounded below by the average degree of a graph $G$, that is, whether there exist a universal constant $c>0$ such that
(12) $\frac{c}{|V|}\sum_{x\in V}d_{x}\leq\nu(G).$
Example 2.19 answers this question negatively.
Example 2.19 (Counterexample to Question 2.18).
A counterexample for the left hand side of the inequality (12) are increasing circles $C_{d}$, since $\nu(C_{d})\to 0$ as $d\to\infty$, while the average degree of these graphs is constant and equal to $2$.
Question 2.20.
Restricting to signatures, the following question arises: Does the smallest eigenvalue $\lambda^{\overline{\sigma}}_{1}(G)$ with anti-balanced signature $\overline{\sigma}\equiv-1$ attain the maximum among all the smallest eigenvalues $\{\lambda^{\sigma}_{1}\}$ corresponding to all signatures $\sigma:E^{or}(G)\to\{\pm 1\}$? A counterexample is given in Example 2.21.
Example 2.21 (Counterexample to Question 2.20).
Let $x_{0},x_{1},x_{2},x_{3}$ be the vertices of the wheel graph $W_{4}$ in the cycle $C_{4}\subset W_{4}$ and $x\in W_{d}$ be the suspension vertex. We choose the following signatures/magnetic potentials:
$(u,v)$ $(x,x_{0})$ $(x,x_{1})$ $(x,x_{2})$ $(x,x_{3})$ $(x_{0},x_{1})$ $(x_{1},x_{2})$ $(x_{2},x_{3})$ $(x_{3},x_{0})$
$\overline{\sigma}(u,v)$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$
$\sigma_{1}(u,v)$ $1$ $1$ $1$ $1$ $-1$ $-1$ $1$ $-1$
$\hat{\sigma}(u,v)$ $-1$ $e^{-i\pi/4}$ $i$ $e^{i\pi/4}$ $1$ $1$ $-1$ $1$
Note that $\overline{\sigma}$ is the anti-balanced signature, while $\sigma_{1}$ is a non-equivalent unbalanced signature. Maple yields the following results:
$\lambda_{1}^{\overline{\sigma}}(W_{4})$ $=1,$
$\lambda_{1}^{\sigma_{1}}(W_{4})$ $=\frac{10}{3}-\frac{\sqrt{19}}{3}\left(\sqrt{3}\cos\left(\frac{\arctan\left(\frac{330\sqrt{426}}{809}\right)}{6}\right)-\sin\left(\frac{\arctan\left(\frac{330\sqrt{426}}{809}\right)}{6}\right)\right)$
$\approx 1.238442816,$
$\lambda_{1}^{\hat{\sigma}}(W_{4})$ $=3-2\cos\left(\frac{\pi}{4}\right)=1+\nu(C_{4})\approx 1.585786438.$
Since $\lambda_{1}^{\hat{\sigma}}(W_{4})\geq\lambda_{1}^{\overline{\sigma}}(W_{4})$, the maximum for $\lambda_{1}^{\sigma}(W_{4})$ amongst all magnetic potentials is not assumed by the anti-balanced signature. Moreover, the unbalanced signature $\sigma_{1}$ provides a strictly larger first magnetic eigenvalue than the anti-balanced signature $\overline{\sigma}$.
3. General results
3.1. Upper combinatorial bounds
We start this section with a combinatorial upper bound of the magneto-spectral height of a graph $G=(V,E)$. To formulate it, we recall that the first Betti number $b_{1}(G)$ of $G$ is given by $|E|-|V|+k$, where $k$ is the number of connected components of $G$. Note that $b_{1}(G)$ is also called the dimension of the cycle space. We have the following three upper combinatorial bounds.
Proposition 3.1.
Let $G=(V,E)$ be a graph. Then we have
$\nu(G)\leq\min\left\{\frac{2|E|}{|V|},\frac{4b_{1}(G)}{|V|},d_{\max}-1\right\},$
where $d_{\max}$ is the maximal vertex of degree of $G$.
Remark 3.2.
All three upper bounds in this proposition are relevant. Note that the upper bound $\frac{2|E|}{|V|}$ is the average vertex degree of the graph $G=(V,E)$. Figure 8 illustrates examples of graphs $G$ for which we have
$\frac{2|E|}{|V|}=\frac{2(2n-1)}{n}<\min\left\{\frac{4b_{1}(G)}{|V|},d_{\max}-1\right\},$
(since $n\geq 6$, $\frac{4b_{1}(G)}{|V|}=4$ and $d_{\max}=n-2$) and
$\frac{4b_{1}(G)}{|V|}=\frac{4(e-n+1)}{n}<\min\left\{\frac{2|E|}{|V|},d_{\max}-1\right\},$
(since $n\geq 5$, $e=|E|=2n-3$, $\frac{2|E|}{|V|}=4-\frac{6}{n}$ and $d_{\max}=n-2$), respectively. Examples of graphs for which we have
$d_{\max}-1<\min\left\{\frac{2|E|}{|V|},\frac{4b_{1}(G)}{|V|}\right\},$
are all $d$-regular connected graphs with $d\geq 3$.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{n-3}$$x_{n-2}$$x_{n-1}$$x_{n}$
Figure 8. If the dashed circles are included and $n\geq 6$, the resulting graph $G$ satisfies $\frac{2|E|}{|V|}<\frac{4b_{1}(G)}{|V|},d_{\max}-1$. If the dashed circles are dropped and $n\geq 5$, the resulting graph $G$ satisfies $\frac{4b_{1}(G)}{|V|}<\frac{2|E|}{|V|},d_{\max}-1$.
Proof of Proposition 3.1.
The proof of $\nu(G)\leq\frac{2|E|}{|V|}$ is a simple trace argument. For $x\in V$, let $d_{x}$ be the vertex degree of $x$. Then we have for every $\sigma:E^{or}(G)\to S^{1}$,
$\lambda_{1}^{\sigma}(G)\leq\frac{1}{|V|}\sum_{j=1}^{|V|}\lambda_{j}^{\sigma}(G)=\frac{1}{|V|}{\rm{Trace}}(\Delta^{\sigma})=\frac{1}{|V|}\sum_{x\in V}d_{x}=\frac{2|E|}{|V|}.$
The upper bound $\nu(G)\leq\frac{4b_{1}(G)}{|V|}$ is a consequence of the more general Theorem 3.3(a) in the next subsection, where we choose $G_{0}=G$, which implies $d_{x}^{out}=0$ for any $x\in G$.
Finally, the upper bound $d_{\max}-1$ is a direct consequence of (10) in Lemma 2.4.
3.2. Upper bounds in terms of subgraphs
In this section, we present the following relations between $\nu(G)$ and properties of subgraphs $G_{0}$.
Theorem 3.3.
Let $G=(V,E)$ be a graph and $G_{0}=(V_{0},E_{0})$ be a non-empty subgraph of $G$. Let $d_{x}^{out}:=|\{y\in V\setminus V_{0}:y\sim x\}|$ be the out-degree of $x\in V_{0}$. Then $\nu(G)$ has the following upper bounds.
  • (a)
    We have
    (13) $\nu(G)\leq\frac{4b_{1}(G_{0})}{|V_{0}|}+\frac{1}{|V_{0}|}\sum_{x\in V_{0}}d_{x}^{out}.$
  • (b)
    We have
    (14) $\nu(G)\leq\nu(G_{0})+\max_{x\in V_{0}}d_{x}^{out}.$
Moreover, if there exists an eigenfunction $f\in C(V_{0},\mathbb{C})$ of the smallest eigenvalue of a maximal magnetic potential of $G_{0}$ with constant modulus $|f|$, then we have
(15) $\nu(G)\leq\nu(G_{0})+\frac{1}{|V_{0}|}\sum_{x\in V_{0}}d_{x}^{out}.$
Note that the upper bound (15) is generally stronger than both (13) and (14), since $\frac{4b_{1}(G_{0})}{|V_{0}|}\geq\nu(G_{0})$ and $\max_{x\in V_{0}}d_{x}^{out}\geq\frac{1}{|V_{0}|}\sum_{x\in V_{0}}d_{x}^{out}$. Moreover, the upper bound (15) holds whenever we choose $G_{0}$ to be a tree or a cycle in $G$. In the special case that $G_{0}$ runs through all edges, we recover the upper bound (10) in Lemma 2.4.
Proof.
In the two statements (a) and (b) of the theorem, we choose different functions and calculate the corresponding Reighley quotients, respectively.
  • (a)
    We first consider the case that $G_{0}$ is connected. Assume $T\subset G_{0}$ is a spanning tree of $G_{0}=(V_{0},E_{0})$. We choose $x_{0}$ as a base point of $T$. Let $\hat{\sigma}:E^{or}(G)\to S^{1}$ be a maximal magnetic potential of $G$. Then we choose the test function $f$ as follows:
    $f(z)=\begin{cases}1,&\text{if $z=x_{0}$,}\\ \hat{\sigma}(x_{0},x_{1})^{-1}\hat{\sigma}(x_{1},x_{2})^{-1}\cdots\hat{\sigma}(x_{n-1},z)^{-1},&\text{if $z\in V_{0}$,}\\ 0,&\text{if $z\notin V_{0}$.}\end{cases},$
    where, in the second case, $x_{0}\sim x_{1}\sim\cdots x_{n-1}\sim z$ is the unique path from $x_{0}$ to $z$ on the spanning tree $T$. Then we have
    $\nu(G)=\lambda^{\hat{\sigma}}_{1}(G)\leq\mathcal{R}^{\hat{\sigma}}(f)$ =$=\frac{\sum_{\{x,y\}\in E}|f(x)-\hat{\sigma}(x,y)f(y)|^{2}}{\sum_{x\in V}|f(x)|^{2}}$
    $=\frac{4(|E_{0}|-|V_{0}|+1)+\sum_{x\in v_{0}}d_{x}^{out}}{|V_{0}|}=\frac{4b_{1}(G_{0})}{|G_{0}|}+\frac{1}{|V_{0}|}\sum_{x\in V_{0}}d_{x}^{out}.$
    In the case that $G_{0}$ has more than one connected component, we choose base points in each of its components and argue accordingly.
  • (b)
    Assume $\hat{\sigma}$ is a maximal magnetic potential on $G$. Denote the restriction of $\hat{\sigma}$ to the subgraph $G_{0}$ by $\hat{\sigma}_{0}$. Let $f$ be an eigenfunction on $G_{0}$ corresponding to the smallest eigenvalue $\lambda^{\hat{\sigma}_{0}}_{1}(G_{0})$ and denote its trivial extension to $G$ by $\tilde{f}$. Then the Reighley quotient of $\tilde{f}$ is given as follows:
    (16) $\nu(G)$ $=\lambda_{1}^{\hat{\sigma}}(G)\leq\mathcal{R}^{\hat{\sigma}}(\tilde{f})$
    =$=\frac{\sum_{\{x,y\}\in E}|\tilde{f}(x)-\hat{\sigma}(x,y)\tilde{f}(y)|^{2}}{\sum_{x\in V}|\tilde{f}(x)|^{2}}$
    (17) =$=\frac{\sum_{\{x,y\}\in E_{0}}|f(x)-\hat{\sigma}_{0}(x,y)f(y)|^{2}}{\sum_{x\in V_{0}}|f(x)|^{2}}+\frac{\sum_{x\in V_{0}}\sum_{y\notin V_{0},x\sim y}|f(x)|^{2}}{\sum_{x\in V_{0}}|f(x)|^{2}}$
    $\leq\lambda^{\hat{\sigma}_{0}}_{1}(G_{0})+\max_{x\in V_{0}}d^{out}_{x}\leq\nu(G_{0})+\max_{x\in V_{0}}d_{x}^{out}.$
    Moreover, in the special case where the norm $|f|$ is constant on $G_{0}$, the second term in (17) becomes the average out-degree, and we obtain the upper bound given in (15).
Corollary 3.4 (Suspension of a graph).
Let $G=(V,E)$ be a graph and $\tilde{G}=(\tilde{V},\tilde{E})$ be its suspension, that is, we add a new vertex $x$ to $G$ and connect it to all vertices in $G$. Then, we have
$\nu(\tilde{G})\leq\nu(G)+1.$
Moreover, in the special case that $G$ is a tree, this is Example 2.10.
Proof.
We choose $G$ as the subgraph of $\tilde{G}$. For each vertex $x\in G$, we have $d_{x}^{out}=1$, and the above inequality follows from Theorem 3.3(b). ∎
3.3. Upper and lower Ramanujan type bounds
A connected $d$-regular graph $G=(V,E)$ is called Ramanujan, if all of its eigenvalues $\lambda\not\in\{0,2d\}$ lie within the (Ramanujan) interval $[d-2\sqrt{d-1},d+2\sqrt{d-1}]$. This interval is the (continuous) $L^{2}$-Laplace spectrum of the $d$-regular tree. The following upper bound is derived in the spirit of the proof of the Alon-Boppana Theorem (see [N91]). Recall that the girth of a graph is the length of a shortest cycle in the graph.
Proposition 3.5.
Let $G=(V,E)$ be a connected graph with $d_{\max}\geq 2$ and $\rm{girth}(G)\geq 2k+1$. Then we have
$\nu(G)\leq d_{\max}-2\sqrt{d_{\max}-1}+\frac{2\sqrt{d_{\max}-1}-1}{k}.$
Proof.
Let $x_{0}\in V$ be a fixed vertex. Then the girth condition implies that the subgraph induced by the vertices of the ball $B_{k}(x_{0})$ of radius $k$ about $x_{0}$ is a tree. For simplicity, we denote the sphere $S_{i}(x_{0})$ of radius $i$ about $x_{0}$ henceforth by $S_{i}$. Let $f:V\to\mathbb{R}$ be defined as follows: $f(x)=F_{i}$ for all $x\in S_{i}(x_{0})$ for $i\leq k$, where
$F_{0}=F_{1}=1,\quad F_{i}=\frac{1}{(\sqrt{d-1})^{i-1}},$
and $f(x)=0$, otherwise. Let $\sigma:E^{or}(G)\to S^{1}$ be an arbitrary magnetic potential. Since this potential can be gauged to become trivial on all edges of $B_{k}(x_{0})$, we can assume, without loss of generality, that $\sigma$ is trivial on these edges, and the Rayleigh quotient of $f$ is given as follows
$\mathcal{R}(f)=\frac{\sum_{i=1}^{k-1}|E(S_{i},S_{i+1})|\cdot(F_{i+1}-F_{i})^{2}+|E(S_{k},S_{k+1})|\cdot F_{k}^{2}}{\sum_{i=0}^{k}|S_{i}|\cdot F_{i}^{2}}.$
The denominator of $\mathcal{R}(f)$ is bounded below by
(18) $\sum_{i=1}^{k}|S_{i}|\cdot F_{i}^{2}=\sum_{i=1}^{k}\frac{|S_{i}|}{(d_{\max}-1)^{i-1}}.$
The first term of the numerator can be estimated from above as follows:
$\sum_{i=1}^{k-1}|E(S_{i},S_{i+1})|\cdot(F_{i+1}-F_{i})^{2}$ $\leq\sum_{i=1}^{k-1}|S_{i}|(d_{\max}-1)\left(\frac{1}{(\sqrt{d_{\max}-1})^{i}}-\frac{1}{(\sqrt{d_{\max}-1})^{i-1}}\right)^{2}$
(19) $=(d_{\max}-2\sqrt{d_{\max}-1})\sum_{i=1}^{k-1}\frac{|S_{i}|}{(d_{\max}-1)^{i-1}}$
The second term of the numerator can be estimated from above as follows:
$|E(S_{k},S_{k+1})|\cdot F_{k}^{2}$ $\leq|S_{k}|(d_{\max}-1)\cdot\frac{1}{(d_{\max}-1)^{k-1}}$
$=\frac{(d_{\max}-2\sqrt{d_{\max}-1})\cdot|S_{k}|}{(d_{\max}-1)^{k-1}}+\frac{(2\sqrt{d_{\max}-1}-1)\cdot|S_{k}|}{(d_{\max}-1)^{k-1}}$
(20) $\leq\frac{(d_{\max}-2\sqrt{d_{\max}-1})\cdot|S_{k}|}{(d_{\max}-1)^{k-1}}+(2\sqrt{d_{\max}-1}-1)\cdot\frac{1}{k}\sum_{i=1}^{k}\frac{|S_{i}|}{(d_{\max}-1)^{i-1}}.$
In the derivation of (20), we applied the inequality
$|S_{i+1}|\leq(d_{\max}-1)|S_{i}|$
recursively and applied averaging.
The total numerator, that is, the sum of (19) and (20), amounts to
$\left(d_{\max}-2\sqrt{d_{\max}-1}+\frac{2\sqrt{d_{\max}-1}-1}{k}\right)\sum_{i=1}^{k}\frac{|S_{i}|}{(d_{\max}-1)^{i-1}}.$
Combining this with the denominator estimate (18) above, we end up with the estimate
$\mathcal{R}^{\sigma}(f)\leq d_{\max}-2\sqrt{d_{\max}-1}+\frac{2\sqrt{d_{\max}-1}-1}{k}.$
It follows from Proposition 3.5 that we have for any sequence $G_{n}$ of connected $d$-regular graphs ($d\geq 3$) with ${\rm{girth}}(G_{n})\to\infty$:
$\limsup_{n\to\infty}\nu(G_{n})\leq d-2\sqrt{d-1}.$
Next we present a lower Ramanujan bound of the magneto-spectral height for regular graphs.
Proposition 3.6.
Let $G=(V,E)$ be a $d$-regular graph with $d\geq 3$. Then we have
$d-2\sqrt{d-1}\leq\nu(G)\leq d-1.$
The upper bound is assumed for the complete graph $K_{d+1}$.
Proof.
Let $G=(V,E)$ be a $d$-regular graph with $d\geq 3$. By Theorem 5.5 in [MSS15], there exists a signature $\sigma:E^{or}\to\{-1,1\}$ such that
$\lambda_{1}^{\sigma}(G)\geq d-2\sqrt{d-1},$
which implies the lower bound. The upper bound follows straightforwardly from (10) in Lemma 2.4. ∎
The previous two propositions imply, in particular, that we have for any infinite family $G_{n}$ of connected, increasing $d$-graphs, $d\geq 3$ with increasing girth, that the sequence $\nu(G_{n})$ is convergent with
$\lim_{n\to\infty}\nu(G_{n})=d-2\sqrt{d-1},$
which is the lower end of the Ramanujan interval.
Question 3.7.
From this result it is natural to ask whether every graph with minimal vertex degree $d_{\min}\geq 3$ satisfies $\nu(G)\geq d_{\min}-2\sqrt{d_{\min}-1}$.
A weaker question is whether for every pair $3\leq d_{\min}\leq d_{\max}$ there exists $\epsilon>0$, only depending on $d_{\min}$ and $d_{\max}$, such that all graphs with
$d_{\min}\leq d_{x}\leq d_{\max}$
satisfy
$\nu(G)>\epsilon.$
This is false for $d_{\min}=2$, since $\nu(C_{n})\to 0$, and true for $d_{\min}=d_{\max}\geq 3$, by Proposition 3.6.
3.4. Lower bounds in terms of diameter
The lower diameter bounds in this subsection and their proofs are inspired by [ChL24, Theorem 4.2] and [ChL24, Theorem 1.1], respectively.
Proposition 3.8.
Let $G=(V,E)$ be a connected graph. Then we have
$\nu(G)\geq\inf_{\sigma:E^{or}(G)\to\{\pm 1\}\atop\text{unbalanced}}\lambda_{1}^{\sigma}(G)\geq\frac{1}{({\rm{diam}(G)+1})|V|}.$
Note that the order of this inequality is sharp, since we have for cycles $\nu(C_{d})\sim\frac{1}{d^{2}}$ and ${\rm{diam}}(C_{d})\sim d$ and $|C_{d}|=d$.
Proof.
Let $\sigma:E^{or}(G)\to\{\pm 1\}$ be an unbalanced magnetic potential and $f:V\to\mathbb{R}$ be an eigenfunction to the eigenvalue $\lambda_{1}^{\sigma}(G)$. Let $x_{0}\in V$ be a vertex satisfying
$|f(x_{0})|\geq f(x)\quad\text{for all $x\in V$. }$
Without loss of generality, we can assume that $f(x_{0})=1$, by rescaling. This implies that
$\sum_{x\in V}|f(x)|^{2}\leq|V|.$
For any path $x_{0}\sim x_{1}\sim\cdots\sim x_{t}$, we have, by Cauchy-Schwarz,
$\sum_{j=0}^{t-1}|f(x_{j})-\sigma(x_{j},x_{j+1})f(x_{j+1})|^{2}=|f(x_{0})-\sigma(x_{0},x_{1})f(x_{1})|^{2}\\ +|\sigma(x_{0},x_{1})f(x_{1})-\sigma(x_{0},x_{1})\sigma(x_{1},x_{2})f(x_{2})|^{2}+\cdots\\ +|\sigma(x_{0},x_{1})\cdots\sigma(x_{t-2},x_{t-1})f(x_{t-1})-\sigma(x_{0},x_{1})\cdots\sigma(x_{t-1},x_{t})f(x_{t})|^{2}\geq\\ \frac{1}{t}|f(x_{0})-\sigma(x_{0},x_{1})\cdots\sigma(x_{t-1},x_{t})f(x_{t})|^{2}.$
Claim: There exists a path $x_{0}\sim x_{1}\sim\cdots\sim x_{s}$ of length $s\leq{\rm{diam}}(G)+1$, such that
$\sum_{j=0}^{s-1}|f(x_{j})-\sigma(x_{j},x_{j+1})f(x_{j+1})|^{2}\geq\frac{1}{{\rm{diam}}(G)+1}.$
The claim is proved as follows: In the case that there exists $x\in V$ with $f(x)=0$, then we choose a path of length $t\leq{\rm{diam}}(G)$ from $x_{0}$ to $x=x_{t}$, and this inequality yields
$\sum_{j=0}^{t-1}|f(x_{j})-\sigma(x_{j},x_{j+1})f(x_{j+1})|^{2}\geq\frac{1}{t}|f(x_{0})|^{2}\geq\frac{1}{{\rm{diam}}(G)}.$
Now we consider the case when $f(x)\neq 0$ for all $x\in V$. Since $\sigma$ is unbalanced, there must exist an edge $x\sim y$ with
$f(x)\sigma(x,y)f(y)<0$
for, otherwise, $\tau(x):=f(x)/|f(x)|$ would be a switching function such that $\sigma^{\tau}$ is the trivial magnetic potential. Now we choose a path of length $t\leq{\rm{diam}}(G)$ from $x_{0}$ to one of $x,y$, which does not contain the other vertex. Without loss of generality the end point of this path, denoted by $x_{t}$, is $x$. If
$\sigma(x_{0},x_{1})\cdots\sigma(x_{t-1},x_{t})f(x_{t})\leq 0,$
then we obtain
$\sum_{j=0}^{t-1}|f(x_{j})-\sigma(x_{j},x_{j+1})f(x_{j+1})|^{2}\geq\frac{1}{t}|f(x_{0})|^{2}\geq\frac{1}{{\rm{diam}}(G)}.$
Otherwise, we conclude that
$\sigma(x_{0},x_{1})\cdots\sigma(x_{t-1},x_{t})\sigma(x_{t},y)f(y)\leq 0,$
and we obtain, by applying the above inequality again for the extended path from $x_{0}$ to $y=x_{t+1}$ via $x=x_{t}$,
$\sum_{j=0}^{t}|f(x_{j})-\sigma(x_{j},x_{j+1})f(x_{j+1})|^{2}\geq\frac{1}{t+1}|f(x_{0})|^{2}\geq\frac{1}{{\rm{diam}}(G)+1}.$
This finishes the proof of the claim.
The claim implies that
$\lambda_{1}^{\sigma}(G)=\mathcal{R}(f)\geq\frac{1}{{\rm{diam}}(G)+1}\cdot\frac{1}{\sum_{x\in V}|f(x)|^{2}}\geq\frac{1}{({\rm{diam}(G)+1})|V|}.$
The next result holds for graphs with non-negative Bakry-Émery curvature. Before we state it, let us briefly introduce this curvature notion. For $f,g:V\to\mathbb{C}$ and $\sigma:E^{or}(G)\to S^{1}$, we consider the following bilinear forms:
$2\Gamma^{\sigma}(f,g)(x)$ $=-\Delta(f\bar{g})(x)+f(x)\cdot\overline{\Delta^{\sigma}g(x)}+\Delta^{\sigma}f(x)\cdot\overline{g(x)}$
$=\sum_{y:y\sim x}\left(f(x)-\sigma(x,y)f(y)\right)\overline{\left(g(x)-\sigma(x,y)g(y)\right)},$
$2\Gamma_{2}^{\sigma}(f,g)(x)$ $=\Delta\Gamma^{\sigma}(f,g)(x)-\Gamma^{\sigma}(f,\Delta^{\sigma}g)(x)-\Gamma^{\sigma}(\Delta^{\sigma}f,g)(x).$
We say that a graph $G$ satisfies the curvature-dimension inequality $CD^{\sigma}(K,n)$ at a vertex $x\in V$, if we have for all functions $f:V\to\mathbb{C}$,
$\Gamma_{2}^{\sigma}(f,f)(x)\geq\frac{1}{n}(\Delta^{\sigma}f(x))^{2}+K\Gamma^{\sigma}(f,f)(x).$
We say that $G$ satisfies $CD^{\sigma}(K,n)$ globally, if it satisfies $CD^{\sigma}(K,n)$ at all vertices. In the special case of the trivial signature $\sigma_{0}$, we use the simplified notation that $G$ satisfies $CD(K,n)$ instead of $CD^{\sigma_{0}}(K,n)$. This curvature notion was originally introduced in [B85, BE85]. The modification for magnetic potentials was given in [LMP19].
The $CD^{\sigma}$-condition at a vertex $x\in V$ is a local condition, and it is completely determined by the combinatorial structure and the magnetic potential on the incomplete $2$-ball of $x$. The incomplete $2$-ball $\mathring{B}_{2}(x)$ is the induced subgraph of the $2$-ball $B_{2}(x)$ with all spherical edges within the $2$-sphere $S_{2}(x)$ removed.
Proposition 3.9.
Let $G=(V,E)$ be a connected, non-bipartite, triangle-free graph satisfying $CD(0,\infty)$. Then we have
$\nu(G)\geq\frac{C}{{\rm{diam}}^{2}(G)}$
with a universal constant $C$.
Proof.
Step 1: In this step, we prove the following: If $G=(V,E)$ is a connected graph satisfying $CD^{\sigma}(0,\infty)$ for an unbalanced signature $\sigma:E^{or}(G)\to\{-1,1\}$, then
$\lambda_{1}^{\sigma}(G)\geq\frac{C}{{\rm{diam}}^{2}(G)}$
for a universal constant $C$. Since $\sigma$ is unbalanced, we know that $\lambda_{1}^{\sigma}=\lambda_{1}^{\sigma}(G)>0$. Let $f:V\to\mathbb{R}$ be an eigenfunction of $\Delta^{\sigma}$ to the eigenvalue $\lambda_{1}^{\sigma}$. Let
$g(x)=2\Gamma^{\sigma}f(x)+\alpha\lambda_{1}^{\sigma}f^{2}(x)$
and $v\in V$ be such that
$g(v)=\max_{x\in V}g(x).$
Then we have
$0\leq\Delta g(v)$ $=2\Delta\Gamma^{\sigma}f(v)+\alpha\lambda_{1}^{\sigma}\Delta f^{2}(v)$
$=2\Delta\Gamma^{\sigma}f(v)+2\alpha\lambda_{1}^{\sigma}(f(v)\Delta^{\sigma}f(v)-\Gamma^{\sigma}f(v)).$
On the other hand, we conclude from the $CD^{\sigma}(0,\infty)$-condition
$4\Gamma_{2}^{\sigma}(f)(v)=-(2\Delta\Gamma^{\sigma}f)(v)+4\Gamma^{\sigma}(f,\Delta^{\sigma}f)(v)\geq 0$
Combining both results yields
00 0 $\leq 4\Gamma^{\sigma}(f,\Delta^{\sigma}f)(v)+2\alpha\lambda_{1}^{\sigma}(f(v)\Delta^{\sigma}f(v)-\Gamma^{\sigma}f(v))$
$=2\left(2\lambda_{1}^{\sigma}\Gamma^{\sigma}(f)(v)+\alpha(\lambda_{1}^{\sigma})^{2}f^{2}(v)-\alpha\lambda_{1}^{\sigma}\Gamma^{\sigma}f(v)\right)$
$=2\left(\alpha(\lambda_{1}^{\sigma})^{2}f^{2}(v)+(2-\alpha)\lambda_{1}^{\sigma}\Gamma^{\sigma}(f)(v)\right).$
Assuming $\alpha>2$, this implies
$\Gamma^{\sigma}f(v)\leq\frac{\alpha}{\alpha-2}\lambda_{1}^{\sigma}f^{2}(v).$
Using this, we obtain for any $x\in V$:
$g(x)\leq g(v)$ $=2\Gamma^{\sigma}f(v)+\alpha\lambda_{1}^{\sigma}f^{2}(v)$
$\leq\frac{2\alpha}{\alpha-2}\lambda_{1}^{\sigma}f^{2}(v)+\alpha\lambda_{1}^{\sigma}f^{2}(v)$
$\leq\frac{\alpha^{2}}{\alpha-2}\lambda_{1}^{\sigma}\max_{z\in V}f^{2}(z).$
This implies that we have for all $x\in V$,
$\sum_{y\sim x}|f(x)-\sigma(x,y)f(y)|^{2}=2\Gamma^{2}f(x)\leq g(x)\leq\frac{\alpha^{2}}{\alpha-2}\lambda_{1}^{\sigma}\max_{z\in V}f^{2}(z).$
Without loss of generality, we can scale $f$ in such a way that
$f(x_{0})=\max_{z\in V}f(z)=1$
and
$f(x)\geq-1\quad\text{for all $x\in V$.}$
The claim in the proof of Proposition 3.8 guarantees the existence of a path $x_{0}\sim x_{1}\sim\cdots\sim x_{s}$ of length $s\leq{\rm{diam}}(G)+1$, satisfying
$\sum_{j=0}^{s-1}|f(x_{j})-\sigma(x_{j},x_{j+1})f(x_{j+1})|^{2}\geq\frac{1}{{\rm{diam}}(G)+1}.$
Using the above estimate, we conclude that
$\frac{1}{{\rm{diam}}(G)+1}\leq\sum_{j=0}^{s-1}\frac{\alpha^{2}}{\alpha-2}\lambda_{1}^{\sigma}\max_{z\in V}f^{2}(z)\leq\frac{\alpha^{2}}{\alpha-2}({\rm{diam}}+1)\lambda_{1}^{\sigma}.$
This finishes the proof of step 1.
Step 2: Assume that $G=(V,E)$ is a connected, non-bipartite triangle-free graph satisfying $CD(0,\infty)$. Let $\overline{\sigma}:E^{or}(G)\to S^{1}$ be the anti-balanced magnetic potential. Since $G$ is not bipartite, $\overline{\sigma}$ is an unbalanced magnetic potential.
Triangle-freeness guarantees that, at every vertex $x\in V$, the magnetic potential on the incomplete $2$-ball $\mathring{B}_{2}(x)$ can be gauged to be trivial. Since the $CD^{\sigma}(K,n)$ condition at a vertex $x\in V$ is determined by the incomplete $2$-ball $\mathring{B}_{2}(x)$ and does not change under gauge transforms, we conclude that $G$ satisfies also the $CD^{\sigma}(0,\infty)$-condition at all vertices. Then the statement of the proposition follows from Step 1. ∎
4. Graph constructions
In this section, we discuss the behaviour of our magneto-spectral invariant under various graph constructions.
4.1. Connecting two non-adjacent vertices
Lemma 4.1.
Let $G=(V,E)$ be a graph and $\tilde{G}=(V,\tilde{E})$ be obtained from $G$ by adding an edge between two distinct vertices of $G$. Then we have
$\nu(\tilde{G})\geq\nu(G).$
This is a straightforward consequence of the Rayleigh quotient, since this graph manipulation adds an additional non-negative term in the numerator of the Rayleight quotient. An immediate consequence of the lemma and the results in Example 2.2 is the following
Corollary 4.2.
Assume that a graph $G=(V,E)$ with $N$ vertices admits a Hamiltonian cycle. Then we have
$\nu(G)\geq\nu(C_{N})=2-2\cos(\pi/N).$
4.2. Adding a bridge between two graphs
It is obvious that we have for two disjoint graphs $G$ and $H$:
(21) $\nu(G\cup H)=\min(\nu(G),\nu(H)).$
Lemma 4.3.
Let $G=(V,E)$ and $H=(V^{\prime},E^{\prime})$ be two disjoint graphs and $\tilde{G}=(\tilde{V},\tilde{E})$ be obtained by connecting $G$ and $H$ by an edge (bridge). Then we have
(22) $\min(\nu(G),\nu(H))\leq\nu(\tilde{G})\leq\max(\nu(G),\nu(H)).$
Proof.
The lower bound of (22) follows directly from (21) and Lemma 4.1.
For the upper bound, we choose $\hat{\sigma}:E^{or}(\tilde{G})\to S^{1}$ to satisfy
$\nu(\tilde{G})=\lambda_{1}^{\tilde{\sigma}}(\tilde{G}).$
Let $x_{0}\in V$ and $z_{0}\in V^{\prime}$ be the bridge in $\tilde{G}$. Then we have for any nonzero function $f:\tilde{V}\to\mathbb{C}$,
$\mathcal{R}^{\hat{\sigma}}(f)=\\ \frac{{\displaystyle{\sum_{\{x,y\}\in E}}}|f(x)-\hat{\sigma}(x,y)f(y)|^{2}+{\displaystyle{\sum_{\{z,w\}\in E^{\prime}}}}|f(z)-\hat{\sigma}(z,w)f(w)|^{2}+|f(x_{0})-\hat{\sigma}(x_{0},z_{0})f(z_{0})|^{2}}{\sum_{x\in V}|f(x)|^{2}+\sum_{z\in V^{\prime}}|f(z)|^{2}}.$
Let $\sigma_{1}$ and $\sigma_{2}$ be the restrictions of $\hat{\sigma}$ to the graphs $G$ and $H$, respectively. Let $f_{1}:V\to\mathbb{C}$ and $f_{2}:V^{\prime}\to\mathbb{C}$ be non-zero functions satisfying
$\mathcal{R}^{\sigma_{1}}(f_{1})=\lambda_{1}^{\sigma_{1}}(G)\quad\text{and}\quad\mathcal{R}^{\sigma_{2}}(f_{2})=\lambda_{1}^{\sigma_{2}}(H).$
In the case that $f_{1}(x_{0})$ and $f_{2}(z_{0})$ are either both zero or both non-zero, we can assume, after rescaling, that $|f_{1}(x_{0})-\hat{\sigma}(x_{0},z_{0})f_{2}(z_{0})|=0$, and that
$f(v)=\begin{cases}f_{1}(v)&\text{if $v\in V$,}\\ f_{2}(v)&\text{if $v\in V^{\prime}$,}\end{cases}$
is non-zero, too. Then we have
$\mathcal{R}^{\hat{\sigma}}(f)=\frac{\sum_{\{x,y\}\in E}|f(x)-\hat{\sigma}(x,y)f(y)|^{2}+\sum_{\{z,w\}\in E^{\prime}}|f(z)-\hat{\sigma}(z,w)f(w)|^{2}}{\sum_{x\in V}|f(x)|^{2}+\sum_{z\in V^{\prime}}|f(z)|^{2}},$
and applying
$\frac{a+b}{c+d}\leq\max\left\{\frac{a}{c},\frac{b}{d}\right\}$
for $a,b,c,d>0$, we conclude that
$\nu(\tilde{G})$ $=\lambda_{1}^{\hat{\sigma}}(\tilde{G})\leq\mathcal{R}^{\hat{\sigma}}(f)$
$\leq\max\left\{\frac{\sum_{\{x,y\}\in E}|f_{1}(x)-\sigma_{1}(x,y)f_{1}(y)|^{2}}{\sum_{x\in V}|f_{1}(x)|^{2}},\frac{\sum_{\{z,w\}\in E^{\prime}}|f_{2}(z)-\sigma_{2}(z,w)f_{2}(w)|^{2}}{\sum_{z\in V^{\prime}}|f_{2}(z)|^{2}}\right\}$
$=\max\left\{\mathcal{R}^{\sigma_{1}}(f_{1}),\mathcal{R}^{\sigma_{2}}(f_{2})\right\}=\max\{\lambda_{1}^{\sigma_{1}}(G)\,\lambda_{1}^{\sigma_{2}}(H)\}\leq\max\{\nu(G),\nu(H)\}.$
It remains to consider the case when precisely one of $f_{1}(x_{0}),f_{2}(z_{0})$ is zero. Assume, without loss of generality, that $f_{1}(x_{0})=0$ and $f_{2}(z_{0})\neq 0$. We set
$f(v)=\begin{cases}f_{1}(v)&\text{if $v\in V$,}\\ 0&\text{if $v\in V^{\prime}$.}\end{cases}$
Note that $f$ is non-zero, and we obtain
$\nu(\tilde{G})$ $=\lambda_{1}^{\hat{\sigma}}(\tilde{G})\leq\mathcal{R}^{\hat{\sigma}}(f)$
$\leq\frac{\sum_{\{x,y\}\in E}|f_{1}(x)-\sigma_{1}(x,y)f_{1}(y)|^{2}}{\sum_{x\in V}|f_{1}(x)|^{2}}$
$=\mathcal{R}^{\sigma_{1}}(f_{1})=\lambda_{1}^{\sigma_{1}}(G)\leq\nu(G).$
A special case of the above lemma is the following, by choosing $H$ to be a single vertex:
Corollary 4.4.
Let $G$ be a graph and $\tilde{G}$ be obtained by adding a dangling edge to $G$. Then we have
$\nu(\tilde{G})\leq\nu(G).$
4.3. Splitting vertices and edges
We first discuss a graph manipulation involving the splitting of a vertex.
Lemma 4.5.
Let $G=(V,E)$ be a graph and $x\in V$ be a vertex of degree $d=d_{1}+d_{2}$. Let $\tilde{G}=(\tilde{V},\tilde{E})$ be the graph obtained by splitting $x$ into two non-adjacent vertices $x_{1},x_{2}$, where $x_{1}$ is adjacent to $d_{1}$ neighbours of $x$ and $x_{2}$ adjacent to the other $d_{2}$ neighbours of $x$. Then we have
$\nu(G)\geq\nu(\tilde{G}).$
Proof.
Note that under this operation, the set of edges is unchanged (with the minor modification that an edge incident to $x$ is now incident to either $x_{1}$ or $x_{2}$). Let $\sigma:E^{or}(G)\to S^{1}$ be a magnetic potential and $f:V\to\mathbb{C}$ be an arbitrary non-zero function. Let $\tilde{\sigma}:E^{or}(\tilde{G})\to S^{1}$ be the corresponding magnetic potential inheriting the directed edge values of $\sigma$ and $\tilde{f}:\tilde{V}\to\mathbb{C}$ be the following function:
$\tilde{f}(z)=\begin{cases}f(z)&\text{if $z\in V\setminus\{x\}$,}\\ f(x)&\text{if $z\in\{x_{1},x_{2}\}$.}\end{cases}$
If numerator and denominator of $\mathcal{R}^{\sigma}_{G}(f)$ is denoted by $A$ and $B$, then we have
$\mathcal{R}^{\tilde{\sigma}}_{\tilde{G}}(\tilde{f})=\frac{A}{B-|f(x)|^{2}+|\tilde{f}(x_{1})|^{2}+|\tilde{f}(x_{2})|^{2}}=\frac{A}{B+|f(x)|^{2}}\leq\frac{A}{B}=\mathcal{R}^{\sigma}_{G}(f).$
This finishes the proof. ∎
Corollary 4.6.
Let $G=(V,E)$ be a graph whose vertices have even degrees. Then we have
$\nu(G)\geq 2-2\cos\left(\frac{\pi}{|E|}\right).$
Moreover, if this inequality holds with equality, then $G$ is the cycle $C_{|E|}$.
Proof.
Applying the above graph manipulation iteratively, we end up with a new graph $\tilde{G}$ with the same number of edges and which is $2$-regular. Any $2$-regular graph is a disjoint union of cycles. Assume $C_{d}$ ($d\leq|E|$) is the largest cycle in this disjoint union. Using (21) and (6), we conclude that
$\nu(G)\geq\nu(\tilde{G})\geq\nu(C_{d})=2-2\cos\left(\frac{\pi}{d}\right)\geq 2-2\cos\left(\frac{\pi}{|E|}\right).$
This shows
$\nu(G)\geq 2-2\cos\left(\frac{\pi}{|E|}\right).$
If $G$ is not the cycle $C_{|E|}$, then we apply these graph manipulations only until we end up with a graph $\tilde{G}$ with all but one vertices of degree $2$ and the remaining vertex $x$ of degree $4$. The connected component of $\tilde{G}$ containing $x$ must be the union of two disjoint cycles glued at the vertex $x$. In the last step, we split $x$ and disconnect this connected component into two disjoint cycles $C_{j}$ and $C_{k}$ with $j,k<|E|$. Using (21), we conclude that
$\nu(G)\geq\nu(\tilde{G})\geq\nu(C_{k})=2-2\cos\left(\frac{\pi}{k}\right)>2-2\cos\left(\frac{\pi}{|E|}\right),$
proving the rigidity statement. ∎
Remark 4.7.
In view of the above construction, we observe the following: If it is true that there exists a universal $\epsilon>0$ such that all graphs $G$ with minimal vertex degree $\geq 3$ satisfy $\nu(G)>\epsilon$, then it is sufficient to prove this statement for all graphs $G$ with minimal vertex degree $3$ and maximal vertex degree $5$.
The next result is concerned with the splitting of an edge.
Lemma 4.8.
Let $G=(V,E)$ be a graph and $\tilde{G}=(\tilde{V},\tilde{E})$ be obtained by splitting an edge in $G$. Then we have
$\nu(\tilde{G})\leq\nu(G).$
More generally, if $\tilde{G}$ is a subdivision of $G$, then we have
$\nu(\tilde{G})\leq\nu(G).$
Proof.
Let $\{x,y\}\in E$ be the edge which is split with an additional vertex $z$, that is, $\tilde{V}=V\cup\{z\}$ and
$\tilde{E}=\left(E\setminus\{\{x,y\}\}\right)\cup\{\{x,z\},\{z,y\}\}.$
Let $\hat{\sigma}$ be a maximal magnetic potential of $\tilde{G}$, that is,
$\nu(\tilde{G})=\lambda_{1}^{\hat{\sigma}}(\tilde{G}).$
Let $\sigma_{1}$ be the magnetic potential on $G$ which agrees with $\hat{\sigma}$ on the edges $E\setminus\{\{x,y\}\}$ and which satisfies
$\sigma_{1}(x,y)=\hat{\sigma}(x,z)\hat{\sigma}(z,y).$
Let $f:V\to\mathbb{C}$ be a non-zero function satisfying
$\lambda_{1}^{\sigma_{1}}(G)=\mathcal{R}^{\sigma_{1}}(f).$
Let $\tilde{f}:\tilde{V}\to\mathbb{C}$ be the function which agrees with $f$ on the vertices in $V$ and which satisfies
$\tilde{f}(z)=\hat{\sigma}(x,z)^{-1}f(x).$
Then we have
$\nu(\tilde{G})$ $\leq\mathcal{R}^{\hat{\sigma}}(\tilde{f})$
$=\frac{|\tilde{f}(x)-\hat{\sigma}(x,z)\tilde{f}(z)|^{2}+|\tilde{f}(z)-\hat{\sigma}(z,y)\tilde{f}(y)|^{2}+\displaystyle{\sum_{\{v,w\}\in E\setminus\{\{x,y\}\}}}|\tilde{f}(v)-\hat{\sigma}(v,w)\tilde{f}(w)|^{2}}{|\tilde{f}(z)|^{2}+\sum_{w\in V}|\tilde{f}(w)|^{2}}$
$\leq\frac{|f(x)-\hat{\sigma}(x,z)\hat{\sigma}(z,y)f(y)|^{2}+\displaystyle{\sum_{\{v,w\}\in E\setminus\{\{x,y\}\}}}|f(v)-\sigma_{1}(v,w)f(w)|^{2}}{\sum_{w\in V}|f(w)|^{2}}$
$=\mathcal{R}^{\sigma_{1}}(f)=\lambda_{1}^{\sigma_{1}}(G)\leq\nu(G).$
4.4. Cartesian products
Lemma 4.9.
Let $G$ and $H$ be two graphs. Then we have
$\nu(G\times H)\geq\nu(G)+\nu(H).$
Proof.
We use the following fact for magnetic potentials $\sigma_{1}$ and $\sigma_{2}$ on $G$ and $H$, respectively, and their Cartesian product magnetic potential $\sigma$:
${\rm{spec}}^{\sigma}(G\times H)=\left\{\lambda+\mu:\lambda\in{\rm{spec}}^{\sigma_{1}}(G),\mu\in{\rm{spec}}^{\sigma_{2}}(H)\right\},$
where ${\rm{spec}}^{\sigma}$ is the multiset of eigenvalues of the magnetic Laplacian $\Delta^{\sigma}$. Here $\sigma:E^{or}(G\times H)\to S^{1}$ is defined as follows:
$\sigma(\{(x,z),(y,z)\})=\sigma_{1}(x,y)\quad\text{for all $x\sim_{G}y$ and $z\in H$}$
and
$\sigma(\{(x,z),(x,w)\})=\sigma_{2}(z,w)\quad\text{for all $x\in G$ and $z\sim_{H}w$.}$
This implies that
$\lambda_{1}^{\sigma}(G\times H)=\lambda_{1}^{\sigma_{1}}(G)+\lambda_{1}^{\sigma_{2}}(H).$
Therefore, we have
$\nu(G\times H)$ $=\sup_{\sigma}\lambda_{1}^{\sigma}(G\times H)$
$\geq\sup_{\sigma_{1},\sigma_{2}}\left(\lambda_{1}^{\sigma_{1}}(G)+\lambda_{1}^{\sigma_{2}}(H)\right)$
$=\sup_{\sigma_{1}}\lambda_{1}^{\sigma_{1}}(G)+\sup_{\sigma_{2}}\lambda_{1}^{\sigma_{2}}(H)=\nu(G)+\nu(H).$
4.5. Suspensions
Regarding suspensions, we have the following result.
Theorem 4.10.
Let $G$ be an arbitrary connected graph and $\tilde{G}$ be its suspension. Then we have
(23) $11\leq\nu(\tilde{G})\leq\nu(G)+$ 1 .
In particular, we have
(24) $\nu(\tilde{T})=1$
for the suspension $\tilde{T}$ of any tree $T$.
Proof.
The upper bound of $\nu(\tilde{G})$ follows directly from Corollary 3.4. Regarding the lower bound for $\nu(\tilde{G})$, it is sufficient to prove it for trees due to Lemma 4.1. The result (24) for the suspension of trees $T$ follows then immediately from (23) and $\nu(T)=0$.
Let us now provide the proof of the lower bound in (23) for a tree $T=(V,E)$ with $V=\{x_{1},\dots,x_{t}\}$. Let $x$ be the suspension vertex of $\tilde{T}$. Any $\sigma:E^{or}(\tilde{T})\to S^{1}$ can be switched to a magnetic potential which is trivial on the edges of the tree and assumes the values $a_{j}=-e^{i\alpha_{j}}$ on the directed edge $(x,x_{j})\in E^{or}(\tilde{T})$. The corresponding Laplacian is represented by the following matrix:
$\Delta^{\sigma}\cong\begin{pmatrix}\begin{array}[]{c|ccccc}t&a_{1}&a_{2}&\cdots&a_{t-1}&a_{t}\\ \hline\cr\bar{a}_{1}&&&&&\\ \bar{a}_{2}&&&&&\\ \vdots&&&\Delta_{T}+{\rm{Id}}&&\\ \bar{a}_{t-1}&&&&&\\ \bar{a}_{t}&&&&&\end{array}\end{pmatrix},$
where $\Delta_{T}$ is the non-normalized standard Laplacian of the tree $T$. Under the assumption that the values $a_{1},\dots,a_{t}\in S^{1}$ satisfy
(25) $\sum_{j=1}^{t}a_{j}=0,$
the vector $(0,1,\dots,1)^{\top}$ is an eigenvector of $\Delta^{\sigma}$ to the eigenvalue $1$ since any constant vector is an eigenvector of $\Delta_{T}$ to the eigenvalue 00. It remains to prove that $1$ is the smallest eigenvalue of $\Delta^{\sigma}$ for a suitable choice of $a_{1},\dots,a_{t}\in S^{1}$ satisfying (25), since we already know from Example 2.10 that $\nu(\tilde{T})\leq 1$. We prove this by showing that we can choose $a_{1},\dots,a_{t}\in S^{1}$ such that
$\Delta^{\sigma}-{\rm{Id}}\cong\begin{pmatrix}\begin{array}[]{c|ccccc}t-1&a_{1}&a_{2}&\cdots&a_{t-1}&a_{t}\\ \hline\cr\bar{a}_{1}&&&&&\\ \bar{a}_{2}&&&&&\\ \vdots&&&\Delta_{T}&&\\ \bar{a}_{t-1}&&&&&\\ \bar{a}_{t}&&&&&\end{array}\end{pmatrix}$
is positive semidefinite. The corresponding quadratic form for the complex variables $z,z_{1},\dots,z_{t}\in\mathbb{C}$ is given by
(26) $Q=(t-1)|z|^{2}+\sum_{j=1}^{t}(a_{j}z\bar{z}_{j}+\bar{a}_{j}\bar{z}z_{j})+\sum_{j=1}^{t}d_{x_{j}}|z_{j}|^{2}-\sum_{\{x_{j},x_{k}\}\in E}(z_{j}\bar{z}_{k}+\bar{z}_{j}z_{k}).$
We consider $T$ to be a rooted tree with a leaf root $x_{1}$ and assume the other vertices $x_{2},\dots,x_{t}$ to be ordered with respect to the distance to the root. Note that every vertex in $T$ has at most one parent and potentially several children. Let $I$ be the set of index pairs $(k,l)$ corresponding to oriented edges $x_{k}\sim x_{l}$ in $T$ with $x_{k}$ the parent of $x_{l}$. Let
$S_{6}=\{\pm 1,\pm\omega,\pm\omega^{2}\}$
be the set of all $6$-th roots of unity with $\omega=e^{2\pi i/3}$. The following claim allows us to find a suitable choice of $a_{1},\dots,a_{t}\in S^{1}$ with the required properties.
Claim: There exist a family of values $\{b_{jk}\}_{(j,k)\in I}$ in $S_{6}$ which satisfies the following: For any vertex $x_{j}\in T$ we have
(27) $a_{j}:=\left(\sum_{k:(j,k)\in I}b_{jk}\right)-b_{j_{0}j}\in S_{6},$
where $x_{j_{0}}\in T$ is the parent of $x_{j}$. Here we use the convention that, for vertices without parents or without children, the corresponding term is chosen to be zero.
$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{6}$$a_{7}$$a_{8}$$b_{12}$$b_{23}$$b_{24}$$b_{35}$$b_{46}$$b_{47}$$b_{48}$
Figure 9. The relation between the values $a_{j}$ and $b_{jk}$. For example, we have $a_{2}=b_{23}+b_{24}-b_{12}$.
Before proving the claim, we now explain how to derive positive semidefiniteness of the matrix $\Delta^{\sigma}-{\rm{Id}}_{t+1}$ with a magnetic potential corresponding to the $a_{j}$’s, derived from the $b_{jk}$’s via (27) (see Figure 9 for a visual illustration of these values). We first need to verify that we have $\sum a_{j}=0$. This can be verified as follows:
$\sum_{j}a_{j}$ $=\left(\sum_{j\in V}\sum_{k:(j,k)\in I}b_{jk}\right)-\left(\sum_{j\in V}b_{j_{0}(j)j}\right)$
$=\left(\sum_{(j,k)\in I}b_{jk}\right)-\left(\sum_{(j,k)\in I}b_{jk}\right)=0$
where $x_{j_{0}(j)}$ is the parent of $x_{j}$. Next we show positive semidefiniteness of the quadratic form $Q$ in (26): We have
00 0 $\leq\sum_{(j,k)\in I}|b_{jk}z+z_{j}-z_{k}|^{2}$
$=\left(\sum_{(j,k)\in I}|b_{jk}|^{2}\right)|z|^{2}+\sum_{j=1}^{t}d_{x_{j}}|z_{j}|^{2}-\sum_{\{x_{j},x_{k}\}\in E}(z_{j}\bar{z}_{k}+\bar{z}_{j}z_{k})$
$\,\,\,\,\,+\sum_{j=1}^{t}\underbrace{\left(\sum_{k:(j,k)\in I}b_{jk}-b_{j_{0}(j)j}\right)}_{=a_{j}}z\bar{z}_{j}+\sum_{j=1}^{t}\underbrace{\left(\sum_{k:(j,k)\in I}\bar{b}_{jk}-\bar{b}_{j_{0}(j)j}\right)}_{=\bar{a}_{j}}\bar{z}z_{j}$
$=Q,$
since $|I|=t-1$ and $|b_{jk}|=1$.
It remains to show the claim, which we prove via Induction. The base case is $T=K_{2}$ (that is, $t=2$). In this case we choose $b_{12}=1$ and obtain $a_{1}=1$, $a_{2}=-1$. Assume the claim holds for any tree of size $t\geq 2$. Let $T$ be a rooted tree with root $x_{1}$ and $t+1$ vertices. Since $x_{1}$ is a leaf, it has precisely one child, namely $x_{2}$. Let $d=d_{x_{2}}$. Then $x_{2}$ has precisely $d-1$ children $x_{3},\dots,x_{d+1}$, and we can consider the rooted subtrees $T_{j}$, $j=3,4,d+1$ of $T$, where $T_{j}$ has the leaf root $x_{2}$ and is connected to $x_{j}$. Hence, the trees $T_{j}$ have only one vertex in common, namely, $x_{2}$. By induction hypothesis, each $T_{j}$ has a system of variables $\{b_{jk}\}$ corresponding to the edges of $T_{j}$, satisfying the claim. It is easy to see that there exist $b,s_{3},\dots,s_{d+1}\in S_{6}$ such that
$\left(\sum_{j=3}^{d+1}s_{j}b_{2j}\right)-b\in S_{6},$
by choosing the factors $s_{j}$ such that two consecutive terms $s_{2l+1}b_{2l+1}$ and $s_{2l+2}b_{2l+2}$ in the above sum cancel each other out and using the fact that $\xi_{6}^{l}+\xi_{6}^{l+2}+\xi_{6}^{l+4}=0$ for the choice of $b$. A suitable family of $\{b_{jk}\}$ for the tree $T$ is now given by $b_{12}=b$ and $b_{kl}=s_{j}b_{kl}$ for edges $\{x_{k},x_{l}\}$ in the subtree $T_{j}$. This leads to $a_{1}=b_{12}\in S_{6}$ and $a_{2}=\sum s_{j}b_{2j}-b\in S_{6}$, completing the induction step. ∎
5. Relation to the spectral gap
The following two examples show that $\nu(G)$ cannot be bounded above by the spectral gap $\lambda_{2}(G)$.
Example 5.1 (The dumbbell).
Let $G_{d}$ be the union of two copies of $K_{d}$, connected by a bridge. Then we have, by Cheegers’ inequality,
$\lambda_{2}(G_{d})\leq 2h(G_{d})\leq 2\frac{1}{|K_{d}|}\to 0,$
by using the division of $G_{d}$ into the two copies of $K_{d}$. That means that we have
$\lambda_{2}(G_{d})\to 0\quad\text{as $d\to\infty$.}$
On the other hand, we have
$\nu(G_{d})=\nu(K_{d})=d-2,$
by Lemma 4.3, that is
$\nu(G_{d})\to\infty\quad\text{as $d\to\infty$.}$
Example 5.2 (Hypercubes).
Let $Q^{d}$ be the $d$-dimensional hypercube. Then we have
$\lambda_{2}(Q^{d})=\lambda_{2}(K_{2})=2.$
On the other hand, we have
$\nu(Q^{2d})=\nu((C_{4})^{d})\geq d\cdot\nu(C_{4}),$
by Lemma 4.9. Since $\nu(C_{4})>0$, we have
$\nu(Q^{2d})\to\infty\quad\text{as $d\to\infty$.}$
Question 5.3.
We do not have any example which shows that $\lambda_{2}(G)$ cannot by bounded above by $\nu(G)$. So an interesting question is whether there exists a universal $C>0$ such that
$\lambda_{2}(G)\leq C\nu(G)?$
In view of this question, Proposition 3.9 can be viewed as some kind of support of this question, since there exists an equivalent lower diameter bound for the spectral gap:
Theorem 5.4 (see [CLY14, Theorem 3.5]).
Let $G=(V,E)$ be a connected graph satisfying $CD(0,\infty)$. Then we have
$\lambda_{2}(G)\geq\frac{C}{{\rm{diam}}^{2}(G)}$
with a universal constant $C$.
Our final result relating $\nu(G)$ to the spectral gap reads as follows. This confirms the question positively in the special case of vertex transitive graphs which are not bipartite.
Proposition 5.5.
Let $G=(V,E)$ be a connected, non-bipartite vertex transitive graph of degree $d$. Then we have
$\nu(G)\geq\frac{C}{d}\lambda_{2}(G)$
with a universal constant $C>0$.
Proof.
Theorem 1.1 in [Sa23] states the following relation for eigenvalues of the normalized Laplacian in connected, non-bipartite vertex transitive graphs $G=(V,E)$ of degree $d$ with $N=|V|$
$2-\tilde{\lambda}_{N}\geq C\frac{\tilde{\lambda}_{2}}{d}$
with a universal constant $C>0$. This implies for the non-normalized Laplacian that
$C\frac{\lambda_{2}(G)}{d}\leq 2d-\lambda_{N}(G)=\lambda_{1}^{\overline{\sigma}}(G)\leq\nu(G),$
where $\overline{\sigma}\equiv-1$ is the anti-balanced magnetic potential. ∎
6. Some open problems
We end this paper with a collection of questions about the magneto-spectral height.
Problem 6.1 (Multiplicity Problem).
Is there an characterization of connected graphs for which there exists a maximal magnetic potential $\hat{\sigma}$ such that $\lambda_{1}^{\hat{\sigma}}(G)$ has higher multiplicity? Trees are not in this class of graphs and also not $K_{3}$ with a dangling edge attached. However, cycles, wheel graphs and complete graphs are in this class of graphs. This question is somehow in the same spirit as the Colin de Verdière Invariant of graphs.
Problem 6.2 (Universal Lower Bound Problem).
Let $d\geq 3$. Do we have $\nu(G)\geq d-2\sqrt{d-1}$ for all graphs with $d_{min}\geq d$? This lower bound is true for $d$-regular graphs and for graphs which can be obtained from a $d$-regular graph by adding edges between existing vertices. A weaker question is the following: Does there exists $\epsilon>0$, such that all graphs with $d_{min}\geq 3$ satisfy $\nu(G)\geq\epsilon$.
Problem 6.3 (Spectral Gap Problem).
Does there exist a constant $C>0$ such that $\lambda_{2}(G)\leq C\nu(G)$ for all graphs or, at least, for large families of graphs. Such a relation holds for all non-bipartite vertex transitive graphs with a constant depending on the vertex degree.
Problem 6.4 (Suspension Problem).
We know that adding a new vertex to a graph $G$ and connecting this vertex to the graph by at least one edge increases the $\nu$-invariant by at most $1$ (note that this is true for adding dangling edges, since then the $\nu$-invariant is non-increasing). Let $G$ be a graph with at least one edge and $\tilde{G}$ be its suspension. Do we have
$\nu(\tilde{G})=\nu(G)+1?$
The statement holds for complete graphs, since $K_{d+1}$ is the suspension of $K_{d}$ and $\nu(K_{d})=d-2$. It also holds for trees $T$ since we have $\nu(T)=0$ and $\nu(\tilde{T})=1$, if $\tilde{T}$ is the suspension of $T$. There is also supporting evidence from the cycles $C_{d}$, $d\geq 3$, since we have $\nu(W_{d})=\nu(C_{d})+1$.
Problem 6.5 (Algebraicity Problem).
Does there exist a graph $G$ for which $\nu(G)$ is transcendental, or is $\nu(G)$ for all graphs $G$ an algebraic number? A related question is about the nature of maximal magnetic potentials $\hat{\sigma}:E^{or}(G)\to S^{1}$. Are there examples of graphs $G=(V,E)$, $x_{0}\in V$, with $\rho_{\hat{\sigma}}(\gamma)=e^{2\pi i\alpha}$ with a non-algebraic $\alpha\in\mathbb{R}$ for at least one $\gamma\in\pi_{1}(G,x_{0})$? Recall that $\rho_{\hat{\sigma}}$ was introduced in (3).
Problem 6.6 (Cardinality problem of maximal potentials).
Do there exist finite graphs $G$ which have infinitely many non gauge-equivalent maximal magnetic potentials?
Problem 6.7 (Asymptotic Triviality Problem).
Does there exist criteria of graph sequences $G_{n}$ satisfying $\nu(G_{n})\to 0$? Obviously, this is the case for sequences which end eventually in trees or forests or for increasing cycles.
Problem 6.8 (Cospectrality Problem).
Subsection 2.4 provides an example of two non-isomorphic cospectral graphs with different magneto-spectral height. Are there also examples of non-isomorphic graphs which are cospectral and have the same magneto-spectral height? This question is similar in spirit as the topic in [CP23] for pairs of Riemannian manifolds, which deals with the question whether higher-dimensional twisted isospectrality of pairs of manifolds over the same base manifold implies already that they are isometric.
Problem 6.9 (Lift Problem).
Does there exist an example of a finite lift $\widehat{G}\to G$ such that $\nu(\widehat{G})>\nu(G)$? While eigenvalues of magnetic Laplacians are preserved under lifts of the magnetic potential and new eigenvalues may occur, the lift allows for magnetic potentials which are not lifts and which may increase the magneto-spectral height. However, we are not aware of any such example.
Problem 6.10 (Average magneto-spectral height Problem).
Instead of taking the supremum over all magnetic potentials in the definition (2) of $\nu(G)$, one could also take the average where all magnetic potentials are chosen uniformly and independently at random. Does there exist a uniform lower bound of this average $\nu_{\rm{av}}(G)$ in terms of $\nu(G)$?
Problem 6.11 (Small Eigenvalues Problem for hyperbolic surfaces).
Do we have for any sequence $S_{n}$ of compact hyperbolic surfaces with ${\rm{inj}}(S_{n})\to\infty$ (where ${\rm{inj}}(S)$ is the injectivity radius of $S$) that lim $\limsup\nu(S_{n})\leq 1/4$? (This would be a smooth counterpart of the Alon-Boppana type result in Subsection 3.3.) Moreover, is it true that we have $\nu(S)\geq 1/4$ for any compact hyperbolic surface? The bound $1/4$ is the bottom of the spectrum of the hyperbolic plane, which is the universal cover of any hyperbolic surface. Moreover, these questions have natural extensions beyond constant curvature and also beyond 2-dimensional surfaces.
Acknowledgement: C. Hu, S. Kamtue and Sh. Liu are grateful for the hospitality of Durham University. C. Hu is supported by the Outstanding PhD Students Overseas Study Support Program of the University of Science and Technology of China. B. Hua is supported by the National Natural Science Foundation of China No. 12371056. S. Kamtue is supported by grants for development of new faculty staff, Ratchadaphiseksomphot Fund, Chulalongkorn University. Sh. Liu is supported by the National Key R & D Program of China 2020YFA0713100 and the National Natural Science Foundation of China No. 12031017 and No. 12431004.